### v1

# LingoQA: Video Question Answering for Autonomous Driving

Abstract
Autonomous driving has long faced a challenge with public acceptance due to the lack of explainability in the decision-making process. Video question-answering (QA) in natural language provides the opportunity for bridging this gap. Nonetheless, evaluating the performance of Video QA models has proved particularly tough due to the absence of comprehensive benchmarks. To fill this gap, we introduce LingoQA [^1] https://github.com/wayveai/LingoQA, a benchmark specifically for autonomous driving Video QA. The LingoQA trainable metric demonstrates a 0.95 Spearman correlation coefficient with human evaluations. We introduce a Video QA dataset of central London consisting of 419k samples that we release with the paper. We establish a baseline vision-language model and run extensive ablation studies to understand its performance.
1 Introduction
Communication plays a pivotal role in naturally fostering trust among individuals. However, establishing trust between users and agents remains a significant challenge within the field of artificial intelligence. Recent studies have indicated that articulating explicit reasoning steps can significantly enhance user confidence [1], in addition to improving the capabilities of Large Language Models (LLMs) [52]. The need for explainability remains critical, particularly in safety-critical domains where technology adoption hinges upon this factor [31].
Consider the domain of end-to-end autonomous driving [11], where the driving policy is often executed through deep neural networks processing camera inputs to generate control commands. Despite their efficiency, these systems have grappled with the persistent challenge of providing transparent and interpretable decisions. Integrating Vision-Language Models (VLMs) into the field of autonomous driving holds the promise of enhancing user trust in these systems. Recent strides in VLMs have solidified transformers as multimodal learners, showcasing remarkable performance in tasks such as Visual Question Answering (VQA) and underscoring their proficiency in acquiring robust representations for complex tasks [14].
Our focus is on vision-only end-to-end autonomous driving, aiming to bridge the gap between data-driven decision-making and user trust. We introduce LingoQA, a benchmark designed for autonomous driving video QA, utilizing a novel dataset comprising more than 419k QA pairs. Distinguished by its free-form approach to questions and answers, this dataset broadens the scope of autonomous driving video QA, encompassing reasoning and action justifications. Additionally, we publish a comprehensive evaluation suite consisting of 1,000 examples. At the core of our benchmark lies a novel evaluation metric based on a learned text classifier called Lingo-Judge, inspired by GPT-Judge used in TruthfulQA [35]. We perform rigorous studies correlating automatic metrics to human preferences and find that Lingo-Judge achieves a 0.950 Spearman and 0.993 Pearson correlation coefficient, surpassing existing automated labelling techniques like METEOR [5], BLEU [42], CIDEr [49], and GPT-4 [41] on our benchmark, while being fast enough for frequent runs during training and development. The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking of autonomous driving explainability.
Equipped with this evaluation toolkit, we conducted a comprehensive empirical study on key components and their ablations in VLMs for autonomous driving. Our findings in Section 5 indicate that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13], on both Action and Scenery datasets. This process involves using 5 video frames over 4 seconds and a late video fusion technique. Our collective work, spanning the LingoQA benchmark, the visual instruction-tuning dataset, and the innovative evaluation metric, aims to propel the domain of explainable autonomous driving, laying a robust foundation for subsequent research and development endeavors. To summarise the main contributions of this paper:
  • LingoQA Benchmark: We introduce a novel benchmark for autonomous driving video QA using a learned text classifier for evaluation. It outperforms existing metrics, including GPT-4, with a Spearman coefficient of 0.950 indicating a strong correlation with human evaluation.
  • LingoQA Dataset: Our 419.9k QA pair dataset stands out with its free-form questions and answers, covering not just perception but also driving reasoning from the drivers directly, broadening the scope of autonomous driving video QA.
  • LingoQA Baseline: Through testing of various video-language components on LingoQA, we find that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13] and a late video fusion technique. We establish a new baseline for this field with an identified model combination. Example outputs from the model are shown in Figure LABEL:fig:lingo_examples.
2 Related work
2.1 Language in Autonomous Driving
Modern autonomous vehicle software relies heavily on artificial intelligence models [6, 20, 21, 23]. This, together with the increased number of such vehicles on the road, poses a fundamental challenge in terms of interpretability in the decision-making process [4]. Understanding why a decision is made is crucial for understanding areas of uncertainty, building trust, enabling effective human-AI collaboration, and ensuring safety [54]. In a survey conducted by Partners for Automated Vehicle Education (PAVE) in 2020 [1], 60% of participants stated that they would trust AVs more if they better understood the underlying process of the system. To establish trust with the general public, the inner workings of systems must be explained in a human-interpretable way, such as through language and visual explanations.
The field of autonomous driving has been embracing the opportunity to make end-to-end driving models more explainable using language, understanding that methods based purely on attention [31] are not sufficient [27]. The early explorations of GPT3.5 [46, 39] and GPT4-V [53] on autonomous driving scenarios show that LLMs/VLMs demonstrate superior performance in scene understanding and causal reasoning compared to existing autonomous systems. Works such as ADAPT [29] and LLM-Driver [10] propose multi-task learning frameworks for jointly predicting language and control outputs. Inspired by progress in large language models [47, 13, 60, 41], vision-language models [50, 7, 51, 57, 58, 59, 12, 3, 32, 41, 36, 16] and multi-modal transformers for robotics [19, 9, 8]. Closely related to our proposed baseline is DriveGPT [56], proposing a multi-modal vision-language-action model that tokenizes videos, as well as text and control actions.
2.2 Evaluation Metrics
Progress has been relatively slow for developing vision-language models for autonomous driving, with only a few works aiming to quantitatively improve upon prior work [30, 29, 56]. A key challenge consists of automated, reproducible evaluation metrics that are highly correlated with human ratings, particularly due to the inherent complexities in assessing natural language. ADAPT [29] reports human feedback in addition to standard natural language generation metrics, while DriveGPT [56] reports ChatGPT ratings. Automated methods such as BLEU [42], METEOR [5], ROUGE [34] show weak alignment with human feedback [49]. CIDEr [49] is also based on n-gram level similarity, as opposed to capturing the correctness of an answer based on its meaning. Newer evaluation metrics using ChatGPT have shown improvement in the area of sentence understanding, while still having limitations, such as providing high scores to elaborate, eloquent, but incorrect sentences [2]. Evaluation based on human feedback is subjective and difficult to reproduce. In this work, we address this challenge by introducing a novel video QA benchmark for autonomous driving that checks for factual correctness and is highly correlated to human correctness ratings on our proposed evaluation dataset.
2.3 Datasets for Autonomous Driving
Recent advances in generative AI have been underpinned by training with increasingly large and diverse internet-scale datasets. Nonetheless, when evaluated zero-shot on autonomous driving datasets, many pre-trained models such as Flamingo [3] and BLIP-2 [32] fall short of expectations. Consequently, there remains a need for high-quality training datasets for autonomous driving, as well as for reliable benchmarks. Autonomous driving datasets have been focused on commentary, as opposed to question-answer pairs [31, 55]. However, we note that action, justification is a relatively easy task for foundation models, that are capable of efficiently compressing large amounts of data [17]. The advantage of VQA is that related, but slightly different questions can be asked to truly probe the validity of the model representations. Further datasets focus on ranking the importance of elements in the scene and explaining the reasoning behind the choice [45, 38]. The datasets available for VQA have been previously constructed around existing object detection datasets [43, 18]. Datasets such as NuScenesQA [43], while useful for testing perception, contain simple language outputs of on average one word per question that do not tackle the more challenging reasoning problem.
Our proposed dataset LingoQA addresses the existing gap in autonomous driving as it contains a diverse set of questions related to driving behaviours and scenery in addition to perception questions related to object presence and positioning. This dataset has the strength of being diverse with respect to the language used while being grounded in human reasoning. Examples of the complex questions and answers existent in the dataset are provided in Appendix A. Moreover, a significant advantage of LingoQA is that the intentions and reasoning behind driving decisions are directly labeled by the drivers themselves. This provides an accurate understanding of driving dynamics.
3 LingoQA Benchmark
In this section, we introduce LingoQA, a benchmark to evaluate video question-answering models for autonomous driving.
3.1 Evaluation Dataset
We collected a small, low-density dataset from in-house human labelers, creating both the questions and the answers associated with the short videos. We labeled a small portion of held-out data on 500 human-generated questions using 20+ different evaluators to obtain our test set. Since answers are subjective and noisy, we labeled them twice, making sure the same evaluator does not receive the same question twice. After that, we manually reviewed the answers for semantic disagreements and mistakes. We relabeled such samples two more times and fixed the disagreements, preferring the semantics of the majority of responders but preserving maximal variety in the responses. Finally, we condensed this into 1k high-quality answers to 500 questions, with two correct but diverse answers per question. The dataset evaluates a range of competencies, including action and justification, attention, description, localisation, identification, counting and anticipation, as shown in Figure 2.
3.2 Evaluation Metric
Evaluating open-ended textual dialogues is a challenging task. Quite often the correct answers are ambiguous, subjective, or even not attainable. The most common language-based metrics for evaluating question-answering models in autonomous driving [56, 29, 31] are BLEU [42], METEOR [5] and CIDEr [49], despite their known limitations, such as relying heavily on the n-gram frequency as opposed to the underlying meaning of the answer. To address these limitations, we set ourselves the challenge to develop an automated, non-visual evaluation method for free-form language answers from vision-language models which checks correctness independent of phrasing against a ground truth answer and which is highly correlated with human ratings.
Pearson Spearman Val Acc. [%] Time [sec]
Lingo-Judge 0.993 0.950 95.0 10.5
GPT-4 with CoT 0.990 0.932 91.2 3016.0
GPT-4 [41] 0.988 0.941 90.6 812.4
BLEU [42] 0.881 0.835 - 0.1
METEOR [5] 0.891 0.876 - 8.0
CIDEr [49] 0.878 0.853 - 0.2
Table 1: Lingo-Judge Performance. Correlation with human ratings, validation accuracy, and time taken to run of our proposed LingoQA evaluation metric compared to previous language-based metrics. All metrics use textual ground truth and have no access to vision information. Further examples are presented in Appendix B.
GPT-4 based evaluation
Inspired by the G-Eval metric [37], we used GPT-4 to evaluate answers on a larger scale. Given a question and answer pair from the test set and a model’s answer, we ask GPT-4 to evaluate whether the model’s answer corresponds to a human’s answer. Notice that it doesn’t make use of any visual input. We experimented with prompts and methods achieving good quality of judgements. We achieved the highest accuracy by employing chain-of-thought prompting where we ask GPT-4 to first come up with an evaluation strategy before grading a model’s answer. However, as shown in Table 1, this leads to increased inference time. Further details are provided in Appendix C. Unfortunately, we found GPT-4 based evaluation impractical to use as a main development and training metric due to the time required to evaluate answers on our relatively small evaluation dataset (from 13min up to 50min for a single evaluation due to the API rate limit).
Lingo-Judge
Given these limitations and inspired by TruthfulQA GPT-Judge [35], we pursued an alternative approach using a learned text classifier, dubbed Lingo-Judge, which estimates the correctness of model answers. We measure the correctness of model predictions as an accuracy using a small transformer-based text classifier that takes in a question, the human’s, and the model’s answer and outputs a probability that the model’s answer is correct. Please note, Lingo-Judge does not receive video input and must rely only on the supporting human’s answers. For every question, we run Lingo-Judge on all combinations of (ground-truth answer, predicted answer) and take the maximum correctness estimate, as shown in Equation 1, where $S$ is the score per sample. We found this recipe yield the best predictive power provided enough diversity of human answers in our evaluation dataset.
$S=\max_{j\in\{0,1\}}F_{Judge}(pred,ground\_truth[j])$ (1)
The architecture of the classifier is a DeBERTa-V3 [22] language model, fine-tuned with LoRA [24]. The classification score is predicted using a linear head on top of the class token output. We fine-tuned the model on a diverse dataset of model predictions from early experiments, where questions and ground truth answers come from our evaluation dataset and the correctness target is labeled by human annotators. On top of this initial dataset, we iteratively improved the classifier using active learning by correcting the wrong predictions of discarded models and adding corrections to the training dataset. On a held-out test set, we find that the binary classification accuracy of the classifier is 95%.
In comparison to metrics such as CIDEr, which provide a system-level performance metric, the classifier provides a probability of correctness for each of the model predictions, meaning that it provides metrics at the sample level. Examples are provided in the Appendix B. This means that 100% classifier accuracy is easy to interpret. The classifier allows us to compute metrics during training, running over our full evaluation dataset in 10 seconds using an A100 GPU.
Correlation to human ratings
We studied empirical correlation of various metrics with human judgments. Several human annotators assigned a scalar score [0, 1] to the inference outputs of 17 different models which can be interpreted as the probability that the response correctly addresses the question [35]. Notably, this process takes several days, highlighting the need for an automated evaluation metric that provides faster development feedback. The final human score of each model is the average of all inference output scores. Further details regarding the methodology for the correlation analysis are in the Appendix D.
The Spearman rank correlation coefficient of our automated metric, Lingo-Judge, with human scores is 0.95, and the Pearson correlation coefficient is 0.993. These values are notably higher compared to other natural language evaluation metrics and GPT-4, as detailed in Table 1. Our analysis demonstrates that Lingo-Judge accurately mirrors human judgments, outperforming existing metrics such as BLEU, METEOR, and CIDEr, as well as GPT-4 with and without chain-of-thoughts prompting. This indicates that Lingo-Judge can effectively serve as a proxy for human labelling, which is particularly significant given the stagnant nature of metrics in autonomous driving since the introduction of the CIDEr metric in 2015. Notably, despite their limitations, prominent models like ADAPT [29] and DriveGPT [56] still use BLEU, METEOR, and CIDEr metrics and report ChatGPT ratings without analyzing their correlation to human preferences. Our work fills this gap by providing a reliable benchmark that better reflects human preferences.
Scenarios QA pairs QA per scenario
Action 24.5k 267.8k $\approx 10.9$
Scenery 3.5k 152.5k $\approx 43.6$
Eval. Dataset 100 1000 $10$
Table 2: Dataset Split. It consists of three different datasets of varying annotation densities. The Action dataset focuses on questions related to driving behaviours, the Scenery dataset focuses on perception capabilities, while the evaluation dataset is designed to probe a range of competencies.
Scenarios Annotations QA Captioning Video length [sec]
Rank2Tell [45] 118 >118>118 20
BDD-OIA [55] 22.9k 35k 5
BDD-X [31] 6.9k 26k 40
NuScenesQA [43] 34k 460k 20
DriveLM [15] 30k 360k 20
LingoQA 28k 419.9k 4
Table 3: Dataset Features. The dataset that we introduce alongside our benchmark consists of questions related to object presence, as well as action, justification, attention, localisation, counting, anticipation, and counterfactuals. In total, it has a similar size to other driving-related datasets such as NuScenesQA, while having a much higher diversity and not being limited to questions related to object positioning.
3.3 Datasets
Figure 2: Dataset Statistics. Dataset split by the number of question-answer pairs for the competencies covered and for the objects referred. One question-answer pair might cover more than one competency or object, hence the total is higher than the size of the datasets. The Action and Scenery datasets have complementary strengths, with one focused more on action-justification competencies and one more on description and localisation.
We created a collection of datasets for explainable driving. The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz. The total size of the dataset is about 10x larger than BDD-X [31], as shown in Table 3. Compared to prior datasets such as NuScenesQA [43], our dataset contains reasoning pairs in addition to object presence, description, and localisation. The answers are also more free-form and more complex, with an average answer length of 17.2 words versus 1.0 words in NuScenesQA. Examples of question answers pairs from LingoQA are shown in Appendix A.
Our labeled autonomous driving training dataset consists of two complementary parts: the action dataset and the scenery dataset.
Action dataset.
This dataset was made from a recorded driving corpus of interesting events where the car’s behavior changes, such as decelerations, accelerations, lane changes, narrow gaps, and turns. Such events were succinctly labeled by driving operators with very short high-level descriptions of the situations and behavioral policies (e.g. “following lane, pedestrian on a zebra crossing, should stop”). Additionally, we added metadata for such events from various perception systems, such as traffic light presence, vehicles and pedestrian visual detectors, weather descriptors, as well as other metadata (speed, steering wheel position, and road type from the map data). Using this data, we developed prompt templates for (1) describing the current action and its justification and (2) a set of example questions and hints about what the answer should mention. Next, we used those prompts with GPT-3.5 to rephrase, answer, and extend the example questions using the provided action description and answer hints. We rebalanced events by bucketing by actions and behavioral policies and sampled up to 500 events from each bucket without replacement, leading to 24,577 video snippets with 167,774 question/answer pairs.
Scenery dataset
The scenery dataset was built to complement the action dataset by focusing on perception-related question in addition to driving behaviours. The dataset was made by densely and thoroughly labelling three 30-minute driving sessions with the ELAN video annotation software[40]. For the entire duration of the driving sessions, we provided short captions in about 15 different categories:
  • Driver’s actions, and their justifications
  • Driver’s attention
  • Observations about relevant vehicles, pedestrians, and other road actors with their visual descriptions
  • Observations about relevant static road elements such as traffic lights, traffic islands, lane and intersection structures
  • Miscellaneous observations about the environment, such as weather, tube stations, and buildings.
Then, for every keyframe every second (1fps), we collect all annotations around this frame and build a textual description containing the driver’s actions and their justifications, the objects requiring the driver’s attention, and the observations. As opposed to the Action dataset, where recommended questions were provided to GPT-3.5 for rephasing, for the Scenery dataset, we asked GPT-4 to generate questions and answers using a set of generic prompts, but also using a prompt with the chain of thought specifically targeting perception questions. This forced GPT-4 to generate many diverse questions and answers. This led to a high-quality diverse dataset with about 43 QA-pairs per video.
Our training dataset covers 9 different competencies: action (what the vehicle is doing), justification (why the action is taken), attention (what should be paid attention to in the current situation), identification (identifying an object given its description), localisation, description, counting, anticipation and reasoning given counterfactuals. The questions also cover a diverse set of objects, such as pedestrians, vehicles, cyclists, buildings, road infrastructure, signs, markings. In Figure 2, we present the number of question and answer pairs for each of the 9 competencies above, as well as for the referred objects, for our two datasets, namely Action and Scenery. The complementary strengths of the datasets are apparent, with one focused on driving behaviours and one on perception tasks.
4 Model Methodology
We propose LingoQA Baseline, a vision language model for autonomous driving based on Vicuna v1.5 [13] with 7B parameters that takes explainability beyond simple image captioning and can answer reasoning questions grounded in video outputs. We train a model that consumes a short video segment and produces answers to autonomous driving-related questions.
4.1 Architecture
The LingoQA Baseline model architecture is based on recent VLMs [32, 36, 19] but enhances them by incorporating a video encoding strategy to process multiple frames from a video snippet, as is shown in Figure 3.
Figure 3: LingoQA Baseline model architecture. We first encode individual frames using CLIP and Q-Former. The Q-Former outputs tokens and we feed the tokens from all frames along with chat history and questions into the LLM, which then predicts an answer.
Vision encoder.
We use CLIP [44], a Vision Transformer (ViT) pre-trained contrastively on image-language pairs, to encode images into features. The inputs to the vision encoder are RGB images from the front camera. We squash the input images to a size of $224\times 224$ as opposed to cropping them in order to keep the full image context. Subsequently, we pass the features through a transformer network, the Querying Transformer (Q-Former), that akin to BLIP-2 [32] acts as a bridge between the vision and language feature spaces. The embeddings are then projected into the large language model (LLM) space using a linear projection layer. We repeat this process for $T=5$ frames of the input video and concatenate the tokens from each image.
Large language model.
We leverage pretrained LLMs to give LingoQA Baseline the ability to answer general questions related to both driving scenes, as well as general knowledge. We use Vicuna v1.5 [13] with 7B parameters built on top of Llama-2 [47]. The language model is auto-regressive and hence can be conditioned on textual inputs, as well as image tokens. The training objective is to predict the next language token in a sequence. We mask all tokens from the training loss that belong to the text prompt, including question and chat history.
4.2 Training Recipe
Our training uses a two-step approach to better utilise video features and improve learning in answering questions based on video data. Through this two-step training, we aim for a better understanding and use of video data in addressing challenges in autonomous driving. During both stages, we train only the attention layers of both CLIP and the large language model, as well as the parameters of the Q-Former and language projection layer, while keeping all other parameters frozen. Further details regarding training parameters are presented in Appendix E.
Ablation Lingo-Judge [%] $\uparrow$ BLEU $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
LingoQA Baseline 60.80 15.00 $18.56$ 65.61
Training recipe Instead of pre-train and fine-tune No fine-tuning $33.60$ $8.33$ $14.33$ $39.16$
No pre-training $56.60$ $13.53$ $17.91$ $57.98$
Fine-tuning dataset Instead of action and scenery Action only $53.80$ $11.65$ $17.68$ $46.50$
Scenery only $55.40$ $13.00$ $18.38$ $55.88$
Frame count Instead of 5 frames Single frame $57.00$ $14.21$ $18.40$ $59.46$
3 frames $59.80$ $14.61$ $18.44$ $62.61$
7 frames $60.60$ $14.46$ 18.61 $61.82$
Video fusion Instead of late-fusion Early-fusion $48.40$ $13.98$ $17.61$ $61.42$
Mid-fusion $59.20$ $14.44$ $18.47$ $63.05$
Language model Instead of Vicuna-1.5-7B[13] OPT-7B [60] $50.00$ $14.98$ $15.99$ $60.08$
Llama-2-7B-Chat [48] $59.20$ $13.52$ $18.43$ $59.87$
Mistral-7B-Instruct [28] $58.00$ $13.80$ $18.33$ $64.21$
Table 4: Empirical Evaluation on LingoQA. Ablation study highlighting the impact of various modifications in training recipes, dataset composition, frame count, video processing techniques and language model.
Stage 1: Pre-training for feature alignment.
In the first stage, we pre-train the model on the GQA and SVIT datasets to align image features with the embedding space of the pretrained LLM. The GQA dataset [25] contains more than 22M questions over 113k images. The recently introduced SVIT dataset [61] contains 4.2M question-answer pairs over 108.1k images. We leverage initial weights from different models to accelerate the training process. We initialise the vision encoder using publicly available weights of OpenCLIP [26], the Q-Former from BLIP2 weights [33], and language model from Vicuna v1.5 [13].
Stage 2: Fine-tuning for video QA.
In the second stage, we fine-tune the model on our video question-answering Action and Scenery datasets described in Section 3.3. During the fine-tuning phase, each sample is composed of 5 frames taken from a 4-second span of video, accompanied by a QA-pair. To facilitate further exploration of autonomous driving QA, we open-source the dataset used to fine-tune LingoQA Baseline.
5 Empirical Evaluation on LingoQA
With the highly modular architecture of VLMs, the question remains what architectural components of the LingoQA Baseline model and dataset composition contribute the most to its performance? We conduct several ablation studies around the architecture and training paradigm described in Section 4. To this end, we investigate variations to the training strategy, training data composition, frame count, video fusion methods, and the use of different large language models. The results are obtained by having each model generate one answer per question and then compare the predicted answer to the two ground truth answers. Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix F.
5.1 Training Strategy
The aim of the training strategy experiments is to understand how much the pre-training and the fine-tuning steps contribute to performance. It becomes apparent that fine-tuning is essential to yield answers relevant to autonomous driving. The model fine-tuned on the LingoQA dataset has nearly double the performance of the dataset that is pre-trained on generic VQA datasets. This shows that while generalised pre-training leads to improvements, task-specific fine-tuning is still required for optimal performance.
5.2 Training Datasets Mixture
Table 4 shows the contributions of our two datasets, Action Dataset and the Scenery Dataset. Both datasets proved influential in improving model performance. We show that fine-tuning on the LingoQA dataset that we open source leads to a considerable improvement compared to general pre-training only.
5.3 Impact of Frame Count
We want to investigate the variation in VQA performance with decreasing and increasing the number of video frames fed into the model. The base model contains 5 frames over a 4-second context. The performance declined when shifting from multi-frame video to a single image representation, which can be explained by the model not getting enough information to answer questions where temporal information is crucial. The performance when increasing the number of frames to 3, 5 and 7 remains relatively consistent. We hypothesise this is due to the lack of effective video fusion and leave it to future work to investigate video-language encoders further,
5.4 Impact of Video Fusion Strategy
Given how crucial temporal context is for scenario understanding, this study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion. The early-fusion method employs average pooling to condense features from the vision encoder prior to their incorporation into the Q-Former, producing a unified visual feature vector for language space projection. In contrast, the mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism. The late-fusion method, where individual frame embeddings from Q-Former output are fed into the LLM, allowing it to resolve temporal relationships. Our findings demonstrate that both mid-fusion and late-fusion are effective methods for incorporating video content into the model. Mid-fusion allows for a greater number of context tokens through the use of a predetermined number of video embeddings. Conversely, late-fusion shows a slightly enhanced performance by providing comprehensive frame information to the LLM.
5.5 Impact of Large Language Model
We investigate the impact that different Large Language Models have on the overall performance of our vision-language model. As shown in Table 4, the best score is achieved by the Vicuna-1.5-7B [13] that our base model uses. In the same model family Llama-2-7B [48] achieves comparable, but slightly lower performance. Mistral-7B [28], despite its promise of superior performance over Llama-2, proved less effective in our fine-tuning task. Notably, OPT-7B [60] demonstrates significantly lower performance, despite having a similar number of parameters. This discrepancy underscores the crucial role of the pretraining phase in the base language model’s effectiveness.
6 Discussion
Limitations of Lingo-Judge.
This work still has a few limitations that we discuss and provide guidelines as to how their effects may be mitigated. The primary limitation of our proposed Lingo-Judge is that we have not studied its ability to generalise to other datasets and to other domains. The classifier has been trained for autonomous driving evaluation and has been shown effective for this purpose. Hence, the classifier can be used for evaluation when paired with the benchmark dataset that we release with the paper but would need to be re-trained if employed on a new dataset. Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset. The same response style is adopted in the LingoQA training sets and the models. We leave generalization to different answering styles to future work. Notably, Lingo-Judge does rate human labels as highly truthful, providing confidence in its ability to rate typical human answers. Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer. We leave it to future work to extend the classifier to predict human preference scores directly instead of factual correctness.
Strengths of Lingo-Judge.
The strength of our contribution comprises proposing a classifier that is highly correlated with human inputs and efficient to run. In conjunction with the evaluation dataset that we propose, it becomes a useful tool for benchmarking vision-language models for autonomous driving on the video question answering task, which has been historically challenging to evaluate in a consistent fashion. With this contribution, autonomous driving research can be accelerated by providing a reliable, efficient, and easy-to-interpret benchmark.
Dataset and model limitations.
One of the primary constraints is that our model operates on relatively short video segments and few frames, limiting the contextual understanding of scenarios. We also do not test for driving decisions and attention mechanisms, focusing on question-answering abilities only. We did not test the scaling in our models and focused on the most practical 7B parameter LLMs only. Our dataset and baseline are limited to information from a single front-facing car camera, excluding additional sensory inputs like LiDAR that could enrich the model’s understanding of the driving environment. Expanding the model to address the short video context, as well as adding action prediction and evaluation to the dataset and the benchmark would result in a more robust and versatile system for autonomous driving.
7 Conclusion
In this paper, we introduced a novel benchmark for Video Question Answering for autonomous driving. The benchmark consists of a evaluation dataset, learned classifier-based metric Lingo-Judge that is highly correlated with human evaluation, a comprehensive high-quality training dataset for autonomous driving. The fast feedback from employing Lingo-Judge facilitates effective exploration in the video QA field. Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of explainability in end-to-end autonomous driving systems. The LingoQA benchmark is openly released to spur further community research, providing a reliable and highly correlated evaluation method to human ratings.
Acknowledgements
This work was possible through the help of many colleagues across Wayve. We would like to thank Elahe Arani for the valuable discussions and feedback on the report. In particular we would like to acknowledge the support from: Anthony Hu, Miguel Sanchez Lohff, Lorenzo Bertoni, Charlie Lyons-Rothbart, Emma Wang, Harriett-Rose Follas, Kyle Esecson, Ben Foxall, Naama Zahavi, Ruben Diaz, Rudi Rankin, Tilly Pielichaty, Sarah Belghiti, Giulio D’Ippolito, Dan Reisman, Alex Persin, Fergal Cotter, Przemyslaw Mazur, Thomas Sajot, Giacomo Gallino, Alex Garcia Mayans, Tim Geypens, Robin Tweedie, Rebecca Hills.
Appendix A LingoQA Dataset Examples
Further examples on the capabilties existent in the training and the evaluation datasets are shown in Figure 4. The scenery dataset contains highly descriptive elements, such as object colours, junction type, construction zones, traffic lights, and the road layout. The action dataset is complementary and focused on driving competencies, such as the impact of traffic lights on driving and interactions with other road agents. The evaluation dataset contains a broad range of questions aimed to test competencies relevant for autonomous driving.
Figure 4: LingoQA dataset examples. From left to right: scenery dataset, action dataset, and evaluation dataset.
Appendix B Lingo-Judge Examples
We present additional qualitative examples from our evaluation dataset in Table 5, alongside predictions from our base model and corresponding metrics for each individual sample. Metrics based on n-gram matching such as CIDEr tend to be error-prone. For example, expressions that have the same meaning, but entirely different words, are marked as not similar at all, such as “None” and “There are no cars.”. Sentences with minor but significant differences are graded as highly similar, despite having opposite meanings, such as “The traffic lights are showing green” and “The traffic lights are showing red”. Lingo-Judge demonstrates robustness against these varied expressions and subtle changes. Lingo-Judge also has limitations, primarily seen when establishing the correctness of the answer would require extra context from the videos. These examples can be seen in Table 6.
We qualitatively compare our classifier to GPT-4 ratings. These examples are shown in Figure 5. In this situation, GPT-4 is misled by the fact that the model answer contains partially correct information. The GPT-4 assessment states that “The student correctly identified the presence of a traffic light” and, despite the colours not being correct, further explains that “and accurately stated its colour”. This highlights some challenges faced by GPT-4 when trying to rate the truthfulness of an answer. Lingo-Judge correctly identifies that the statements described by the model are false.
Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
How many pedestrians are crossing the road? Zero pedestrians There are no pedestrians crossing the road. 23.18 5 5 0.96 True
What is the road speed limit? 20 mph - it is written on the road The road speed limit is 20 mph. 97.91 4 3 0.95 True
How many cars are driving in your direction? None There are no cars driving in my direction. 0.11 4 3 0.96 True
Which vehicle should you follow if any? The motorcyclist. If any, I should follow the motorcycle ahead. 0.42 4 5 0.95 True
What is the current action and its justification? Answer in the form “action, justification” Slow down, there is a stationary van infront of us I am decelerating because of the stationary truck ahead. 34.79 5 5 0.96 True
What is the current action and its justification? Answer in the form “action, justification” Stop, Red light I am stopping because the traffic lights to go straight are red. 18.62 5 5 0.95 True
How many cyclists can you see? I can see 3 cyclists I can see two cyclists. 150.29 1 2 0.05 False
What color are the traffic lights showing? The traffic lights are showing green The traffic lights are showing red. 329.36 0 1 0.05 False
What action are you taking with respect to the cyclist? Overtaking them on the right and keeping the speed I am overtaking the cyclist on the left. 349.52 2 2 0.10 False
In which direction is the bus driving? The bus is driving in the opposite direction The bus is driving in the oncoming direction. 404.65 4 5 0.31 False
Are there any parked car on the side of the road? Yes, there are two cars parked on the right of the road No, there are no parked cars on either side of the road. 142.40 0 0 0.05 False
Is acceleration necessary in this situation? If so, provide the reason. No. We should decelerate in this situation because there is a vehicle stopping ahead of us. No, acceleration is not necessary in this situation as I am already driving at the speed limit. 177.96 3 3 0.31 False
Table 5: Qualitative comparison of metrics. Questions and labels from our evaluation dataset along with exemplary predictions and their corresponding sample-level metrics. Notable is that the classifier (here “L-J”) can pick up slight but crucial differences between ground truth and prediction to correctly identify a wrong prediction, such as “The traffic lights are green” and “The traffic lights are red”, in which case CIDEr still shows high agreement. Green color indicates agreement with human judgement, while orange and red show disagreement.
Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
Is any vehicle stopped at the zebra crossing? There is no zebra crossing. No, there are no vehicles stopped at the zebra crossing. 106.19 4 2 0.93 True
What elements are capturing your attention and what is the reason behind it? I am making sure there is no vehicle on the left lane I am making my lane change to. I am also looking at the traffic light to make sure it stays green while I am crossing the intersection. I am paying attention to the bus lane on the left because I need to maintain a safe distance from it. 0.91 3 1 0.57 True
What is the current action and its justification? Answer in the form ”action, justification” We are taking a left turn, because we are at a t-junction. Then we keep a safe distance from the cyclists in front. I am decelerating to keep a safe distance from the cyclist ahead of me. 46.45 2 2 0.32 False
Table 6: Failure Cases of Lingo-Judge. Examples where Lingo-Judge makes a wrong judgement about the correctness of the model prediction. Green color indicates agreement with human judgement, while orange and red show disagreement.
Figure 5: Classifier examples. Examples of Lingo-Judge outputs compared to GPT-4.
Appendix C GPT-4 Grading
In this section we provide an overview of the implementation details for the evaluation method using GPT-4 with and without Chain-of-Thought (CoT) [52] prompting.
GPT-4 with CoT. In order to evaluate a model’s answer with GPT-4 and CoT prompting, we first provide GPT-4 with the question and one or more valid answers for the questions, and ask it to come up with a strategy to evaluate new answers to this question. We then provide GPT-4 with the model’s answer and ask it to evaluate the answer using the strategy it proposed in the previous step. Finally, we ask GPT-4 to give the model a grade between 0 and 5, where 5 means the answer is perfect. The prompt used is shown in Figure 6.
GPT-4 without CoT. When evaluating model outputs without CoT prompting, we provide GPT-4 with the question, one or more valid answers for the questions, and the model predictions and we directly ask GPT-4 to give the model a grade between 0 and 5, without the intermeidate reasoning steps. The prompt used is shown in Figure 7.
We emit concurrent requests to our Azure’s GPT-4 deployment in order to max-out the limit of 40k tokens per minute. GPT-4 without CoT prompting required more than 13 minutes to perform the evaluation, and GPT-4 with CoT prompting requires more than 50 minutes.
Figure 6: GPT-4 with Chain of Thought (CoT) prompting. First, GPT-4 is provided with the question and ground truth answers, and asked to come up with a strategy for testing the answer. Second, GPT-4 is provided with the model answer and is prompted to evaluate the accuracy of the response based on the previously defined strategy. Finally, GPT-4 is asked to provide a grade for the student.
Figure 7: GPT-4 without Chain of Thought (CoT) prompting. GPT-4 is provided with a prompt that contains the question, the ground truth answers, and the model response, and is requested to directly provide a grade for the student.
Appendix D Lingo-Judge Correlation Study
We show that Lingo-Judge exhibits a higher correlation to human judgment than commonly-used language-based metrics, and than GPT-4. To do so, we computed the scores of 15 different models and 2 groups of human labellers on the questions in our evaluation dataset using Lingo-Judge, GPT-4, BLEU4, METEOR and CIDEr. These scores are reported in Table 7.
We then computed the Pearson and Spearman correlation coefficients between these metrics and the human evaluation. The Pearson correlation coefficient measures the strength of the linear correlation between the human evaluation and a metric score, while the Spearman rank correlation coefficient measures the monotonic relationship between the human evaluation and the metric. The higher the Spearman coefficient, the better a metric is at ranking answers in the same order as our human evaluators. To compute the confidence intervals, we use the Fisher transformation with a 95% confidence level.
In Figure 8, the metric scores are plotted against the human evaluators’ grades (from 0 to 1). In red is the least-squares regression of the linear relationship between the metric and the human-assigned grades. Figure 9 shows the value of both correlation coefficients for each of the 5 metrics, as well as their confidence interval bounds. We note than not only does Lingo-Judge provided higher correlation, it also provides tighter confidence intervals than the other metrics.
Lingo-Judge [%] $\uparrow$ BLEU $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$ GPT-4 $\uparrow$ Human $\uparrow$
Models Model A $59.6$ $15.45$ $18.36$ $66.32$ $3.23$ $0.571$
Model B $59.6$ $15.16$ $18.84$ $65.11$ $3.16$ $0.564$
Model C $57.4$ $14.87$ $18.52$ $65.49$ $3.08$ $0.563$
Model D $58.2$ $14.51$ $18.59$ $66.02$ $3.15$ $0.559$
Model E $59.0$ $14.42$ $18.58$ $66.95$ $3.14$ $0.553$
Model F $58.0$ $14.82$ $18.89$ $65.43$ $3.11$ $0.552$
Model G $54.8$ $14.41$ $17.86$ $64.67$ $2.98$ $0.529$
Model H $50.0$ $13.29$ $17.44$ $59.87$ $2.88$ $0.520$
Model I $53.0$ $14.63$ $17.98$ $64.45$ $2.96$ $0.510$
Model J $52.6$ $12.17$ $17.59$ $50.45$ $3.00$ $0.509$
Model K $53.0$ $13.20$ $18.03$ $54.90$ $3.04$ $0.500$
Model L $51.2$ $14.69$ $17.83$ $64.51$ $2.91$ $0.485$
Model M $43.2$ $13.76$ $17.37$ $60.36$ $2.67$ $0.371$
Model N $35.8$ $13.18$ $15.67$ $56.07$ $2.41$ $0.361$
Model O $33.6$ $8.33$ $14.33$ $39.16$ $2.07$ $0.279$
Humans Human labellers group A $96.6$ $81.04$ $52.92$ $361.77$ $4.68$ $0.934$
Human labellers group B $91.2$ $61.72$ $42.57$ $267.87$ $4.3$ $0.894$
Table 7: Correlation study metrics. Metrics from different models on our evaluation dataset used in the correlation study in Table 1. For reference, we also present metrics for answers provided by human labellers. “Human” is the average of inference output scores in range $\left[0,1\right]$ where 0 is worst and 1 is best, as described in section 3.2.
Figure 8: Correlation trends. Correlation trends of the average grade of models compared to the average human-grades, for different metrics.
Figure 9: Correlation coefficients. Correlation coefficients of the average grade of different models vs. the average human-grades, for different metrics.
Appendix E Training Parameters
In this sections we present further details on the training parameters used for the LingoQA Baseline. The training process consists of a pre-training stage, and a fine-tuning stage. Table 8 shows the parameters for pre-training and fine-tuning respectively. The datasets are sampled with equal weight for both pre-training and fine-tuning. The overall training time was  20h for pre-training and  5h for fine-tuning on an NVIDIA A100 8GPU 80GB machine.
Parameter Pre-training Fine-tuning
Precision bf16 bf16
Warm-up steps 1000 1000
Maximum steps 100000 10000
Batch size 6 8
Gradient acc. steps 1 1
Learning rate $5*10^{-5}$ $5*10^{-5}$
Learning rate scheduler cosine cosine
Weight decay 0.1 0.1
Table 8: Training parameters. This table shows the training parameters utilised for the pre-training and for the fine-tuning stages respectively.
Appendix F LingoQA Baseline Examples
Figure 10: Examples of model outputs on the LingoQA benchmark. We compare the baseline with a model that has not been fine-tuned on the LingoQA dataset, a model fine-tuned on the action dataset only, and a model fine-tuned on the scenery dataset only. This shows qualitatively how the baseline can handle both action justification as well as descriptive tasks by combining the strengths of both datasets.
We qualitatively showcase the impact of our proposed LingoQA dataset. Figure 10 compares three models: a model that is not fine-tuned on any LingoQA datasets, one that is fine-tuned on the action dataset only, one on the scenery dataset only, and the baseline that is trained with both. Two questions are asked, one focused on perception only, and one focused on action justification. The action only model performs well at answering action-related questions, but not perception. The scenery only model performs well at perception tasks, but not action justification. The baseline exhibits good performance on both.

### v2

@@
 Abstract
-Autonomous driving has long faced a challenge with public acceptance due to the lack of explainability in the decision-making process.
-Video question-answering (QA) in natural language provides the opportunity for bridging this gap.
-Nonetheless, evaluating the performance of Video QA models has proved particularly tough due to the absence of comprehensive benchmarks.
-To fill this gap, we introduce LingoQA [^1] https://github.com/wayveai/LingoQA, a benchmark specifically for autonomous driving Video QA.
-The LingoQA trainable metric demonstrates a 0.95 Spearman correlation coefficient with human evaluations.
-We introduce a Video QA dataset of central London consisting of 419k samples that we release with the paper.
+We introduce LingoQA, a novel dataset and benchmark for video question answering in autonomous driving.
+The dataset contains 28K unique scenarios and 419K annotations.
+Evaluating state-of-the-art vision-language models on our benchmark shows that their performance is below human capabilities, with GPT-4V responding truthfully to 56.67% of the questions compared to 93.4% for humans.
+For evaluation, in addition to conducting a human study, we propose a truthfulness classifier, called Lingo-Judge, that achieves a 0.95 Spearman correlation coefficient to human evaluations, surpassing existing techniques like METEOR, BLEU, CIDEr, and GPT-4.
 We establish a baseline vision-language model and run extensive ablation studies to understand its performance.
+We release our dataset and benchmark [^1] https://github.com/wayveai/LingoQA and hope that it will provide a thorough evaluation platform for future vision-language models in autonomous driving.
+{strip}
+Figure 1: LingoQA is a comprehensive benchmark for Video Question Answering in autonomous driving.
+Our baseline vision-language model on this benchmark, can answer questions related to driving reasoning, object recognition, action justification, and scene description.
 1 Introduction
@@
 Recent studies have indicated that articulating explicit reasoning steps can significantly enhance user confidence [1], in addition to improving the capabilities of Large Language Models (LLMs) [52].
-The need for explainability remains critical, particularly in safety-critical domains where technology adoption hinges upon this factor [31].
+The need for textual justifications remains critical, particularly in safety-critical domains where technology adoption hinges upon this factor [29].
 Consider the domain of end-to-end autonomous driving [11], where the driving policy is often executed through deep neural networks processing camera inputs to generate control commands.
-Despite their efficiency, these systems have grappled with the persistent challenge of providing transparent and interpretable decisions.
+Recent strides in VLMs have solidified transformers as multimodal learners, showcasing remarkable performance in tasks such as Visual Question Answering (VQA) and underscoring their proficiency in acquiring robust representations for complex tasks [14].
 Integrating Vision-Language Models (VLMs) into the field of autonomous driving holds the promise of enhancing user trust in these systems.
-Recent strides in VLMs have solidified transformers as multimodal learners, showcasing remarkable performance in tasks such as Visual Question Answering (VQA) and underscoring their proficiency in acquiring robust representations for complex tasks [14].
 Our focus is on vision-only end-to-end autonomous driving, aiming to bridge the gap between data-driven decision-making and user trust.
@@
 We perform rigorous studies correlating automatic metrics to human preferences and find that Lingo-Judge achieves a 0.950 Spearman and 0.993 Pearson correlation coefficient, surpassing existing automated labelling techniques like METEOR [5], BLEU [40], CIDEr [49], and GPT-4 [39] on our benchmark, while being fast enough for frequent runs during training and development.
-The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking of autonomous driving explainability.
+The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking video question answering in autonomous driving.
 Equipped with this evaluation toolkit, we conducted a comprehensive empirical study on key components and their ablations in VLMs for autonomous driving.
@@
 This process involves using 5 video frames over 4 seconds and a late video fusion technique.
-Our collective work, spanning the LingoQA benchmark, the visual instruction-tuning dataset, and the innovative evaluation metric, aims to propel the domain of explainable autonomous driving, laying a robust foundation for subsequent research and development endeavors.
+Our collective work, spanning the LingoQA benchmark, the visual instruction-tuning dataset, and the innovative evaluation metric, aims to propel the domain of language-augmented autonomous driving, laying a robust foundation for subsequent research and development endeavors.
 To summarise the main contributions of this paper:
@@
 We establish a new baseline for this field with an identified model combination.
-Example outputs from the model are shown in Figure LABEL:fig:lingo_examples.
+Example outputs from the model are shown in Figure LingoQA: Video Question Answering for Autonomous Driving.
 2 Related Work
@@
 In a survey conducted by Partners for Automated Vehicle Education (PAVE) in 2020 [1], 60% of participants stated that they would trust AVs more if they better understood the underlying process of the system.
-To establish trust with the general public, the inner workings of systems must be explained in a human-interpretable way, such as through language and visual explanations.
-The field of autonomous driving has been embracing the opportunity to make end-to-end driving models more explainable using language, understanding that methods based purely on attention [31] are not sufficient [27].
+To establish trust with the general public, the systems must be explained in a human-interpretable way, such as through language and visual explanations [4].
+The field of autonomous driving has been embracing the opportunity to make driving models more trustworthy for their users using visual attention methods [25] or textual explanations [29].
 The early explorations of GPT3.5 [45, 37] and GPT4-V [53] on autonomous driving scenarios show that LLMs/VLMs demonstrate superior performance in scene understanding and causal reasoning compared to existing autonomous systems.
 Works such as ADAPT [27] and LLM-Driver [10] propose multi-task learning frameworks for jointly predicting language and control outputs.
-Inspired by progress in large language models [47, 13, 60, 41], vision-language models [50, 7, 51, 57, 58, 59, 12, 3, 32, 41, 36, 16] and multi-modal transformers for robotics [19, 9, 8].
+Inspired by progress in large language models [47, 13, 60, 39], vision-language models [50, 7, 51, 57, 58, 59, 12, 3, 30, 39, 35, 15] and multi-modal transformers for robotics [17, 9, 8] our work incorporates language to autonomous driving.
 Closely related to our proposed baseline is DriveGPT [56], proposing a multi-modal vision-language-action model that tokenizes videos, as well as text and control actions.
 2.2 Evaluation Metrics
-Progress has been relatively slow for developing vision-language models for autonomous driving, with only a few works aiming to quantitatively improve upon prior work [30, 29, 56].
+Progress has been relatively slow in developing vision-language models for autonomous driving, with only a few works aiming to quantitatively improve upon prior work [28, 27, 56].
 A key challenge consists of automated, reproducible evaluation metrics that are highly correlated with human ratings, particularly due to the inherent complexities in assessing natural language.
@@
 Recent advances in generative AI have been underpinned by training with increasingly large and diverse internet-scale datasets.
-Nonetheless, when evaluated zero-shot on autonomous driving datasets, many pre-trained models such as Flamingo [3] and BLIP-2 [32] fall short of expectations.
-Consequently, there remains a need for high-quality training datasets for autonomous driving, as well as for reliable benchmarks.
-Autonomous driving datasets have been focused on commentary, as opposed to question-answer pairs [31, 55].
-However, we note that action, justification is a relatively easy task for foundation models, that are capable of efficiently compressing large amounts of data [17].
-The advantage of VQA is that related, but slightly different questions can be asked to truly probe the validity of the model representations.
-Further datasets focus on ranking the importance of elements in the scene and explaining the reasoning behind the choice [45, 38].
-The datasets available for VQA have been previously constructed around existing object detection datasets [43, 18].
-Datasets such as NuScenesQA [43], while useful for testing perception, contain simple language outputs of on average one word per question that do not tackle the more challenging reasoning problem.
+[3] [30] This has brought into light the need for evaluation benchmarks and datasets that focus not only on specific tasks, but on reasoning areas, such as descriptive and predictive reasoning.
+[41] Prior works, such as the CausalVQA benchmark [32] and the Perception Test [41], a comprehensive benchmark for vision-language foundation models, probe the validity of the model representations through question answering.
+Autonomous driving datasets have been focused on commentary [29, 55] or constructed around existing object detection datasets [42, 16].
+Datasets such as NuScenesQA [42], contain simple language outputs of on average one word per question that do not tackle the more challenging reasoning problem.
 Our proposed dataset LingoQA addresses the existing gap in autonomous driving as it contains a diverse set of questions related to driving behaviours and scenery in addition to perception questions related to object presence and positioning.
+The evaluation dataset probes areas such as description, counting, localisation, anticipation, attention, and action justification.
 This dataset has the strength of being diverse with respect to the language used while being grounded in human reasoning.
 Examples of the complex questions and answers existent in the dataset are provided in Appendix A.
-Moreover, a significant advantage of LingoQA is that the intentions and reasoning behind driving decisions are directly labeled by the drivers themselves.
-This provides an accurate understanding of driving dynamics.
 3 LingoQA Benchmark
@@
 Given a question and answer pair from the test set and a model’s answer, we ask GPT-4 to evaluate whether the model’s answer corresponds to a human’s answer.
-Notice that it doesn’t make use of any visual input.
+Notice that it does not make use of any visual input.
 We experimented with prompts and methods achieving good quality of judgements.
@@
 For every question, we run Lingo-Judge on all combinations of (ground-truth answer, predicted answer) and take the maximum correctness estimate, as shown in Equation 1, where $S$ is the score per sample.
-We found this recipe yield the best predictive power provided enough diversity of human answers in our evaluation dataset.
-$S=\max_{j\in\{0,1\}}F_{Judge}(pred,ground\_truth[j])$ (1)
+We found that this recipe yields the best predictive power provided enough diversity of human answers in our evaluation dataset.
+$S=\max_{j\in\{0,1\}}F_{\mathrm{Judge}}(\mathrm{prediction},\mathrm{ground\_truth}[j])$ (1)
 The architecture of the classifier is a DeBERTa-V3 [20] language model, fine-tuned with LoRA [22].
@@
 The Spearman rank correlation coefficient of our automated metric, Lingo-Judge, with human scores is 0.95, and the Pearson correlation coefficient is 0.993.
-These values are notably higher compared to other natural language evaluation metrics and GPT-4, as detailed in Table 1.
-Our analysis demonstrates that Lingo-Judge accurately mirrors human judgments, outperforming existing metrics such as BLEU, METEOR, and CIDEr, as well as GPT-4 with and without chain-of-thoughts prompting.
+These values are considerably higher compared to other natural language evaluation metrics and GPT-4, as detailed in Table 1.
+Our analysis demonstrates that Lingo-Judge accurately mirrors human judgments, outperforming existing metrics such as BLEU, METEOR, and CIDEr, as well as GPT-4 with and without chain-of-thought prompting.
 This indicates that Lingo-Judge can effectively serve as a proxy for human labelling, which is particularly significant given the stagnant nature of metrics in autonomous driving since the introduction of the CIDEr metric in 2015.
@@
 Scenarios Annotations QA Captioning Video length [sec]
-Rank2Tell [45] 118 >118>118 20
+Rank2Tell [44] 118 $>118$ 20
 BDD-OIA [55] 22.9k 35k 5
@@
 NuScenesQA [42] 34k 460k 20
-DriveLM [15] 30k 360k 20
+DriveLM [46] 30k 443k 20
 LingoQA 28k 419.9k 4
@@
 The Action and Scenery datasets have complementary strengths, with one focused more on action-justification competencies and one more on description and localisation.
-We created a collection of datasets for explainable driving.
+We created a collection of datasets for bringing language to autonomous driving.
 The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz.
@@
 Scenery dataset
-The scenery dataset was built to complement the action dataset by focusing on perception-related question in addition to driving behaviours.
+The scenery dataset was built to complement the action dataset by focusing on perception-related questions in addition to driving behaviours.
 The dataset was made by densely and thoroughly labelling three 30-minute driving sessions with the ELAN video annotation software[38].
@@
   • Observations about relevant vehicles, pedestrians, and other road actors with their visual descriptions
-  • Observations about relevant static road elements such as traffic lights, traffic islands, lane and intersection structures
+  • Observations about relevant road elements such as traffic lights, traffic islands, lane and intersection structures
   • Miscellaneous observations about the environment, such as weather, tube stations, and buildings.
@@
 4 Model Methodology
-We propose LingoQA Baseline, a vision language model for autonomous driving based on Vicuna v1.5 [13] with 7B parameters that takes explainability beyond simple image captioning and can answer reasoning questions grounded in video outputs.
+We propose LingoQA Baseline, a vision language model for autonomous driving based on Vicuna v1.5 [13] with 7B parameters that can answer reasoning questions grounded in video outputs.
 We train a model that consumes a short video segment and produces answers to autonomous driving-related questions.
 4.1 Architecture
-The LingoQA Baseline model architecture is based on recent VLMs [32, 36, 19] but enhances them by incorporating a video encoding strategy to process multiple frames from a video snippet, as is shown in Figure 3.
+The LingoQA Baseline model architecture is based on recent VLMs [30, 35, 17] but enhances them by incorporating a video encoding strategy to process multiple frames from a video snippet, as shown in Figure 3.
 Figure 3: LingoQA Baseline model architecture.
@@
 4.2 Training Recipe
-Our training uses a two-step approach to better utilise video features and improve learning in answering questions based on video data.
-Through this two-step training, we aim for a better understanding and use of video data in addressing challenges in autonomous driving.
-During both stages, we train only the attention layers of both CLIP and the large language model, as well as the parameters of the Q-Former and language projection layer, while keeping all other parameters frozen.
-Further details regarding training parameters are presented in Appendix E.
+Our training uses a two-step approach to better utilise video features and improve learning when answering questions based on video data.
+Through this two-step training, we aim to better understand and use video data.
+In the first stage, we train the self-attention layers for the LLM and the vision encoder (QKV), together with all the Q-Former parameters and the linear language projection layer.
+In the second stage, we fine-tune the same parameters as in the previous stage, keeping the vision encoder frozen.
+Further details regarding training parameters are presented in Appendix F.
 Ablation Lingo-Judge [%] $\uparrow$ BLEU $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
@@
 The results are obtained by having each model generate one answer per question and then compare the predicted answer to the two ground truth answers.
-Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix F.
+Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix G.
 5.1 Training Strategy
 The aim of the training strategy experiments is to understand how much the pre-training and the fine-tuning steps contribute to performance.
-It becomes apparent that fine-tuning is essential to yield answers relevant to autonomous driving.
-The model fine-tuned on the LingoQA dataset has nearly double the performance of the dataset that is pre-trained on generic VQA datasets.
-This shows that while generalised pre-training leads to improvements, task-specific fine-tuning is still required for optimal performance.
+It becomes apparent that fine-tuning leads to improved answers relevant to autonomous driving.
+The model fine-tuned on the LingoQA dataset has double the performance of the dataset that is pre-trained on generic VQA datasets.
 5.2 Training Datasets Mixture
-Table 4 shows the contributions of our two datasets, Action Dataset and the Scenery Dataset.
+Table 4.2 shows the contributions of our two datasets, Action Dataset and the Scenery Dataset.
 Both datasets proved influential in improving model performance.
@@
 The performance declined when shifting from multi-frame video to a single image representation, which can be explained by the model not getting enough information to answer questions where temporal information is crucial.
-The performance when increasing the number of frames to 3, 5 and 7 remains relatively consistent.
-We hypothesise this is due to the lack of effective video fusion and leave it to future work to investigate video-language encoders further,
+The performance when increasing the number of frames to 3, 5, and 7 remains relatively consistent.
+We hypothesize that this is due to the lack of effective video fusion, hence we highlight the opportunity for improved video encoding methods.
 5.4 Impact of Video Fusion Strategy
-Given how crucial temporal context is for scenario understanding, this study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion.
+Given how crucial the temporal context is for scenario understanding, this study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion.
 The early-fusion method employs average pooling to condense features from the vision encoder prior to their incorporation into the Q-Former, producing a unified visual feature vector for language space projection.
 In contrast, the mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism.
-The late-fusion method, where individual frame embeddings from Q-Former output are fed into the LLM, allowing it to resolve temporal relationships.
+The late-fusion method, where individual frame embeddings from Q-Former output are fed into the LLM, allows it to resolve temporal relationships.
 Our findings demonstrate that both mid-fusion and late-fusion are effective methods for incorporating video content into the model.
@@
 We investigate the impact that different Large Language Models have on the overall performance of our vision-language model.
-As shown in Table 4, the best score is achieved by the Vicuna-1.5-7B [13] that our base model uses.
-In the same model family Llama-2-7B [48] achieves comparable, but slightly lower performance.
+As shown in Table 4.2, the best score is achieved by Vicuna-1.5-7B [13] that our base model uses.
+In the same family of models, Llama-2-7B [48] achieves comparable, but slightly lower performance.
 Mistral-7B [26], despite its promise of superior performance over Llama-2, proved less effective in our fine-tuning task.
@@
 This discrepancy underscores the crucial role of the pretraining phase in the base language model’s effectiveness.
+5.6 Evaluation of SOTA Vision-Language Models
+To demonstrate the relevance of the newly proposed benchmark, we evaluate a series of SOTA vision-language models and compare them to human performance.
+The models shown in Table 5 are evaluated zero-shot.
+The best performing models achieve an accuracy as rated by humans of 56.67%, which is below human performance at 93.4%.
+This highlights the challenging nature of our proposed benchmark.
+Table 5: Evaluating vision-language models zero-shot. The performance of existing vision-language models is far from human capability, highlighting the challenging nature of the benchmark.
+Human Lingo-J GPT-4 BLEU METEOR CIDEr
+Human 93.4 96.6 4.68 81.04 52.92 361.77
+GPT-4V 56.67 59.6 3.19 6.30 12.35 42.82
+LLaVA 38.97 49.4 2.45 4.23 8.38 38.39
+FUYU 17.69 45.4 2.28 1.90 13.00 12.04
 6 Discussion
-Limitations of Lingo-Judge.
-This work still has a few limitations that we discuss and provide guidelines as to how their effects may be mitigated.
-The primary limitation of our proposed Lingo-Judge is that we have not studied its ability to generalise to other datasets and to other domains.
-The classifier has been trained for autonomous driving evaluation and has been shown effective for this purpose.
-Hence, the classifier can be used for evaluation when paired with the benchmark dataset that we release with the paper but would need to be re-trained if employed on a new dataset.
-Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset.
-The same response style is adopted in the LingoQA training sets and the models.
-We leave generalization to different answering styles to future work.
-Notably, Lingo-Judge does rate human labels as highly truthful, providing confidence in its ability to rate typical human answers.
-Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer.
-We leave it to future work to extend the classifier to predict human preference scores directly instead of factual correctness.
 Strengths of Lingo-Judge.
@@
 With this contribution, autonomous driving research can be accelerated by providing a reliable, efficient, and easy-to-interpret benchmark.
+Limitations of Lingo-Judge.
+This work still has a few limitations that we discuss and provide guidelines as to how their effects may be mitigated.
+The classifier has been trained for autonomous driving evaluation and has been shown effective for this purpose.
+Hence, the classifier can be used for evaluation when paired with the benchmark dataset that we release with the paper.
+Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset.
+The same response style is adopted in the LingoQA training sets and the models.
+Further details regarding generalisation to response styles is studied in Appendix E.
+Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer.
 Dataset and model limitations.
@@
 Our dataset and baseline are limited to information from a single front-facing car camera, excluding additional sensory inputs like LiDAR that could enrich the model’s understanding of the driving environment.
-Expanding the model to address the short video context, as well as adding action prediction and evaluation to the dataset and the benchmark would result in a more robust and versatile system for autonomous driving.
+Expanding the model to address the short video context, as well as adding action prediction and evaluation to the dataset and the benchmark would result in a more versatile system for autonomous driving.
 7 Conclusion
@@
 The fast feedback from employing Lingo-Judge facilitates effective exploration in the video QA field.
-Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of explainability in end-to-end autonomous driving systems.
+Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of end-to-end autonomous driving systems.
 The LingoQA benchmark is openly released to spur further community research, providing a reliable and highly correlated evaluation method to human ratings.
@@
 The evaluation dataset contains a broad range of questions aimed to test competencies relevant for autonomous driving.
+Further examples from the evaluation benchmark are also included in the overview video accompanying the submission.
+Notably, any personal identifiable information, such as faces and plate ID’s, has been annonymised in the dataset.
 Figure 4: LingoQA dataset examples. From left to right: scenery dataset, action dataset, and evaluation dataset.
@@
 Lingo-Judge correctly identifies that the statements described by the model are false.
-Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
-How many pedestrians are crossing the road? Zero pedestrians There are no pedestrians crossing the road. 23.18 5 5 0.96 True
-What is the road speed limit? 20 mph - it is written on the road The road speed limit is 20 mph. 97.91 4 3 0.95 True
-How many cars are driving in your direction? None There are no cars driving in my direction. 0.11 4 3 0.96 True
-Which vehicle should you follow if any? The motorcyclist. If any, I should follow the motorcycle ahead. 0.42 4 5 0.95 True
-What is the current action and its justification?
-Answer in the form “action, justification” Slow down, there is a stationary van infront of us I am decelerating because of the stationary truck ahead.
+Question
+Label
+Prediction
+CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
+How many pedestrians are crossing the road?
+Zero pedestrians
+There are no pedestrians crossing the road.
 23.18 5 5 0.96 True
-What is the current action and its justification? Answer in the form “action, justification” Stop, Red light I am stopping because the traffic lights to go straight are red. 18.62 5 5 0.95 True
-How many cyclists can you see? I can see 3 cyclists I can see two cyclists. 150.29 1 2 0.05 False
-What color are the traffic lights showing? The traffic lights are showing green The traffic lights are showing red. 329.36 0 1 0.05 False
-What action are you taking with respect to the cyclist? Overtaking them on the right and keeping the speed I am overtaking the cyclist on the left. 349.52 2 2 0.10 False
-In which direction is the bus driving? The bus is driving in the opposite direction The bus is driving in the oncoming direction. 404.65 4 5 0.31 False
-Are there any parked car on the side of the road? Yes, there are two cars parked on the right of the road No, there are no parked cars on either side of the road. 142.40 0 0 0.05 False
-Is acceleration necessary in this situation?
-If so, provide the reason.
-No.
-We should decelerate in this situation because there is a vehicle stopping ahead of us.
+What is the road speed limit?
+20 mph - it is written on the road
+The road speed limit is 20 mph.
+97.91 4 3 0.95 True
+How many cars are driving in your direction?
+None
+There are no cars driving in my direction.
+0.11 4 3 0.96 True
+Which vehicle should you follow if any?
+The motorcyclist.
+If any, I should follow the motorcycle ahead.
+0.42 4 5 0.95 True
+What is the current action and its justification? Answer in the form “action, justification”
+Slow down, there is a stationary van infront of us
+I am decelerating because of the stationary truck ahead.
+34.79 5 5 0.96 True
+What is the current action and its justification? Answer in the form “action, justification”
+Stop, Red light
+I am stopping because the traffic lights to go straight are red.
+18.62 5 5 0.95 True
+How many cyclists can you see?
+I can see 3 cyclists
+I can see two cyclists.
+150.29 1 2 0.05 False
+What color are the traffic lights showing?
+The traffic lights are showing green
+The traffic lights are showing red.
+329.36 0 1 0.05 False
+What action are you taking with respect to the cyclist?
+Overtaking them on the right and keeping the speed
+I am overtaking the cyclist on the left.
+349.52 2 2 0.10 False
+In which direction is the bus driving?
+The bus is driving in the opposite direction
+The bus is driving in the oncoming direction.
+404.65 4 5 0.31 False
+Are there any parked car on the side of the road?
+Yes, there are two cars parked on the right of the road
+No, there are no parked cars on either side of the road.
+142.40 0 0 0.05 False
+Is acceleration necessary in this situation? If so, provide the reason.
+No. We should decelerate in this situation because there is a vehicle stopping ahead of us.
 No, acceleration is not necessary in this situation as I am already driving at the speed limit.
@@
 Green color indicates agreement with human judgement, while orange and red show disagreement.
-Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
-Is any vehicle stopped at the zebra crossing? There is no zebra crossing. No, there are no vehicles stopped at the zebra crossing. 106.19 4 2 0.93 True
+Question
+Label
+Prediction
+CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
+Is any vehicle stopped at the zebra crossing?
+There is no zebra crossing.
+No, there are no vehicles stopped at the zebra crossing.
+106.19 4 2 0.93 True
 What elements are capturing your attention and what is the reason behind it?
-I am making sure there is no vehicle on the left lane I am making my lane change to.
-I am also looking at the traffic light to make sure it stays green while I am crossing the intersection.
+I am making sure there is no vehicle on the left lane I am making my lane change to. I am also looking at the traffic light to make sure it stays green while I am crossing the intersection.
 I am paying attention to the bus lane on the left because I need to maintain a safe distance from it.
 0.91 3 1 0.57 True
-What is the current action and its justification?
-Answer in the form ”action, justification” We are taking a left turn, because we are at a t-junction.
-Then we keep a safe distance from the cyclists in front.
+What is the current action and its justification? Answer in the form ”action, justification”
+We are taking a left turn, because we are at a t-junction. Then we keep a safe distance from the cyclists in front.
 I am decelerating to keep a safe distance from the cyclist ahead of me.
@@
 Figure 9 shows the value of both correlation coefficients for each of the 5 metrics, as well as their confidence interval bounds.
-We note than not only does Lingo-Judge provided higher correlation, it also provides tighter confidence intervals than the other metrics.
+We note that not only does Lingo-Judge provide higher correlation, it also provides tighter confidence intervals than the other metrics.
+Appendix E Lingo-Judge Generalisation
+To investigate the generalisation abilities of the Lingo-Judge, we examine the performance of the model on a range of answer styles.
+In particular, we evaluate vision-language models with varying architectures namely GPT4-V, LLaVA and FUYU.
+Table 8 shows the performance of the Lingo-Judge as measured by the validation accuracy.
+This highlights that the model performs the best on short, concise answers, akin to those of humans.
+Table 8: Robustness to response styles. Investigation into the impact of the response style on validation accuracy.
+Lingo-J Val. Acc.
+LingoQA 89.50
+GPT-4V
+few-shot (FS) 83.27
+concise (C) 81.63
+unprompted (U) 83.06
+incorrect (I) 89.59
+LLaVA
+concise (C) 81.43
+unprompted (U) 78.12
+FUYU 64.89
 Lingo-Judge [%] $\uparrow$ BLEU $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$ GPT-4 $\uparrow$ Human $\uparrow$
@@
 Figure 9: Correlation coefficients. Correlation coefficients of the average grade of different models vs. the average human-grades, for different metrics.
-Appendix E Training Parameters
+Appendix F Training Parameters
 In this sections we present further details on the training parameters used for the LingoQA Baseline.
@@
 Table 10: Training parameters. This table shows the training parameters utilised for the pre-training and for the fine-tuning stages respectively.
-Appendix F LingoQA Baseline Examples
+Appendix G LingoQA Baseline Examples
 Figure 10: Examples of model outputs on the LingoQA benchmark.

### v4

@@
-LingoQA: Video Question Answering for Autonomous Driving
+https://wayve.ai/science/
 Abstract
-We introduce LingoQA, a novel dataset and benchmark for video question answering in autonomous driving.
-The dataset contains 28K unique scenarios and 419K annotations.
-Evaluating state-of-the-art vision-language models on our benchmark shows that their performance is below human capabilities, with GPT-4V responding truthfully to 56.67% of the questions compared to 93.4% for humans.
-For evaluation, in addition to conducting a human study, we propose a truthfulness classifier, called Lingo-Judge, that achieves a 0.95 Spearman correlation coefficient to human evaluations, surpassing existing techniques like METEOR, BLEU, CIDEr, and GPT-4.
+We introduce LingoQA, a novel dataset and benchmark for visual question answering in autonomous driving.
+The dataset contains 28K unique short video scenarios, and 419K annotations.
+Evaluating state-of-the-art vision-language models on our benchmark shows that their performance is below human capabilities, with GPT-4V responding truthfully to 59.6% of the questions compared to 96.6% for humans.
+For evaluation, we propose a truthfulness classifier, called Lingo-Judge, that achieves a 0.95 Spearman correlation coefficient to human evaluations, surpassing existing techniques like METEOR, BLEU, CIDEr, and GPT-4.
 We establish a baseline vision-language model and run extensive ablation studies to understand its performance.
-We release our dataset and benchmark [^1] https://github.com/wayveai/LingoQA and hope that it will provide a thorough evaluation platform for future vision-language models in autonomous driving.
-{strip}
-Figure 1: LingoQA is a comprehensive benchmark for Video Question Answering in autonomous driving.
-Our baseline vision-language model on this benchmark, can answer questions related to driving reasoning, object recognition, action justification, and scene description.
+We release our dataset and benchmark [^1] https://github.com/wayveai/LingoQA as an evaluation platform for vision-language models in autonomous driving.
+Keywords:
+Autonomous Driving Question Answering Vision-language
+*Equal contributions
+Figure 1: Examples from the LingoQA benchmark designed to evaluate whether models can describe the elements in the scene (description), predict what future action should be taken (action) and why (justification) and explain what they are paying attention to (attention).
 1 Introduction
@@
 Our focus is on vision-only end-to-end autonomous driving, aiming to bridge the gap between data-driven decision-making and user trust.
-We introduce LingoQA, a benchmark designed for autonomous driving video QA, utilizing a novel dataset comprising more than 419k QA pairs.
-Distinguished by its free-form approach to questions and answers, this dataset broadens the scope of autonomous driving video QA, encompassing reasoning and action justifications.
+We introduce LingoQA, a benchmark designed for autonomous driving VQA, utilizing a novel dataset comprising more than 419k QA pairs.
+Distinguished by its free-form approach to questions and answers, this dataset broadens the scope of autonomous driving QA, encompassing reasoning and action justifications.
 Additionally, we publish a comprehensive evaluation suite consisting of 1,000 examples.
@@
 We perform rigorous studies correlating automatic metrics to human preferences and find that Lingo-Judge achieves a 0.950 Spearman and 0.993 Pearson correlation coefficient, surpassing existing automated labelling techniques like METEOR [5], BLEU [39], CIDEr [49], and GPT-4 [38] on our benchmark, while being fast enough for frequent runs during training and development.
-The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking video question answering in autonomous driving.
+The evaluation code and the weights for the classifier will be released with the paper to support robust benchmarking of visual question answering in autonomous driving.
 Equipped with this evaluation toolkit, we conducted a comprehensive empirical study on key components and their ablations in VLMs for autonomous driving.
@@
 To summarise the main contributions of this paper:
-  • LingoQA Benchmark: We introduce a novel benchmark for autonomous driving video QA using a learned text classifier for evaluation.
+  • LingoQA Benchmark: We introduce a novel benchmark for autonomous driving VQA using a learned text classifier for evaluation.
 It outperforms existing metrics, including GPT-4, with a Spearman coefficient of 0.950 indicating a strong correlation with human evaluation.
-  • LingoQA Dataset: Our 419.9k QA pair dataset stands out with its free-form questions and answers, covering not just perception but also driving reasoning from the drivers directly, broadening the scope of autonomous driving video QA.
-  • LingoQA Baseline: Through testing of various video-language components on LingoQA, we find that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13] and a late video fusion technique.
+  • LingoQA Dataset: Our 419.9k QA pair dataset stands out with its free-form questions and answers, covering not just perception but also driving reasoning from the drivers directly, broadening the scope of autonomous driving datasets.
+  • LingoQA Baseline: Through testing of various vision-language components on LingoQA, we find that the most effective approach involves partially fine-tuning the attention layers of our vision-language model equipped with Vicuna-1.5-7B [13] and a late video fusion technique.
 We establish a new baseline for this field with an identified model combination.
-Example outputs from the model are shown in Figure LingoQA: Video Question Answering for Autonomous Driving.
+Example outputs from the model are shown in Figure 1.
 2 Related Work
@@
 Evaluation based on human feedback is subjective and difficult to reproduce.
-In this work, we address this challenge by introducing a novel video QA benchmark for autonomous driving that checks for factual correctness and is highly correlated to human correctness ratings on our proposed evaluation dataset.
+In this work, we address this challenge by introducing a novel VQA benchmark for autonomous driving that checks for factual correctness and is highly correlated to human correctness ratings on our proposed evaluation dataset.
 2.3 Datasets for Autonomous Driving
@@
 This dataset has the strength of being diverse with respect to the language used while being grounded in human reasoning.
-Examples of the complex questions and answers existent in the dataset are provided in Appendix A.
+Examples of the complex questions and answers existent in the dataset are provided in Appendix 0.A.
 3 LingoQA Benchmark
-In this section, we introduce LingoQA, a benchmark to evaluate video question-answering models for autonomous driving.
-3.1 Evaluation Dataset
-We collected a small, low-density dataset from in-house human labelers, creating both the questions and the answers associated with the short videos.
-We labeled a small portion of held-out data on 500 human-generated questions using 20+ different evaluators to obtain our test set.
-Since answers are subjective and noisy, we labeled them twice, making sure the same evaluator does not receive the same question twice.
-After that, we manually reviewed the answers for semantic disagreements and mistakes.
-We relabeled such samples two more times and fixed the disagreements, preferring the semantics of the majority of responders but preserving maximal variety in the responses.
-Finally, we condensed this into 1k high-quality answers to 500 questions, with two correct but diverse answers per question.
-The dataset evaluates a range of competencies, including action and justification, attention, description, localisation, identification, counting and anticipation, as shown in Figure 2.
+In this section, we introduce LingoQA, a benchmark to evaluate autonomous driving reasoning through video question-answering.
+The benchmark consists of an automated evaluation metric and the corresponding datasets for evaluation and fine-tuning.
 3.1 Evaluation Metric
@@
 To address these limitations, we set ourselves the challenge to develop an automated, non-visual evaluation method for free-form language answers from vision-language models which checks correctness independent of phrasing against a ground truth answer and which is highly correlated with human ratings.
+N-Gram Matching Metrics.
+Common language-based metrics such as BLEU [39], METEOR [5] and CIDEr [49] are still the most common metrics used for evaluating captioning in autonomous driving, despite their known limitations, such as relying heavily on the n-gram frequency as opposed to the underlying meaning of the answer.
+The BLEU [39] metric computes the n-gram based precision of the candidate sequence compared to the reference sequence.
+Similarly, ROUGE [33] computes the recall-based n-gram, while METEOR [5] computes the F-measure.
+CIDEr [49] is based on using TF-IDF to provide a lower weight to terms that are commonly reported in the corpus.
+Table 1: Lingo-Judge performance.
+Correlation with human ratings, validation accuracy, and time taken to run of our proposed LingoQA evaluation metric compared to previous language-based metrics.
+All metrics use textual ground truth and have no access to vision information.
+Further examples are presented in Appendix 0.B.
 Pearson Spearman Val Acc. [%] Time [sec]
@@
 CIDEr [49] 0.878 0.853 - 0.2
-Table 1: Lingo-Judge Performance.
-Correlation with human ratings, validation accuracy, and time taken to run of our proposed LingoQA evaluation metric compared to previous language-based metrics.
-All metrics use textual ground truth and have no access to vision information.
-Further examples are presented in Appendix B.
-GPT-4 based evaluation
+GPT-4 based evaluation.
 Inspired by the G-Eval metric [36], we used GPT-4 to evaluate answers on a larger scale.
@@
 However, as shown in Table 1, this leads to increased inference time.
-Further details are provided in Appendix C.
+Further details are provided in Appendix 0.C.
 Unfortunately, we found GPT-4 based evaluation impractical to use as a main development and training metric due to the time required to evaluate answers on our relatively small evaluation dataset (from 13min up to 50min for a single evaluation due to the API rate limit).
-Lingo-Judge
+Lingo-Judge.
 Given these limitations and inspired by TruthfulQA GPT-Judge [34], we pursued an alternative approach using a learned text classifier, dubbed Lingo-Judge, which estimates the correctness of model answers.
@@
 In comparison to metrics such as CIDEr, which provide a system-level performance metric, the classifier provides a probability of correctness for each of the model predictions, meaning that it provides metrics at the sample level.
-Examples are provided in the Appendix B.
+Examples are provided in the Appendix 0.B.
 This means that 100% classifier accuracy is easy to interpret.
 The classifier allows us to compute metrics during training, running over our full evaluation dataset in 10 seconds using an A100 GPU.
-Correlation to human ratings
+Correlation to human ratings.
 We studied empirical correlation of various metrics with human judgments.
@@
 The final human score of each model is the average of all inference output scores.
-Further details regarding the methodology for the correlation analysis are in the Appendix D.
+Further details regarding the methodology for the correlation analysis are in the Appendix 0.D.
 The Spearman rank correlation coefficient of our automated metric, Lingo-Judge, with human scores is 0.95, and the Pearson correlation coefficient is 0.993.
@@
 Our work fills this gap by providing a reliable benchmark that better reflects human preferences.
-Scenarios QA pairs QA per scenario
-Action 24.5k 267.8k $\approx 10.9$
-Scenery 3.5k 152.5k $\approx 43.6$
-Eval. Dataset 100 1000 $10$
-Table 2: Dataset Split.
-It consists of three different datasets of varying annotation densities.
-The Action dataset focuses on questions related to driving behaviours, the Scenery dataset focuses on perception capabilities, while the evaluation dataset is designed to probe a range of competencies.
+3.2 Datasets
+We created a collection of datasets for bringing language to autonomous driving.
+The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz.
+The total size of the dataset is about 10x larger than BDD-X [29], as shown in Table 2.
+Compared to prior datasets such as NuScenesQA [42], our dataset contains reasoning pairs in addition to object presence, description, and localisation.
+The answers are also free-form and more complex, with an average answer length of 17.2 words versus 1.0 words in NuScenesQA.
+Examples of question answers pairs from LingoQA are shown in Appendix 0.A.
+Table 2: Dataset comparison.
+The dataset that we introduce has a similar size to other driving-related while having a much higher diversity.
+A scenario refers to a unique short video sequence, that may have multiple annotations attached to it.
 Scenarios Annotations QA Captioning Video length [sec]
@@
 LingoQA 28k 419.9k 4
-Table 3: Dataset Features.
-The dataset that we introduce alongside our benchmark consists of questions related to object presence, as well as action, justification, attention, localisation, counting, anticipation, and counterfactuals.
-In total, it has a similar size to other driving-related datasets such as NuScenesQA, while having a much higher diversity and not being limited to questions related to object positioning.
-3.3 Datasets
-Figure 2: Dataset Statistics.
-Dataset split by the number of question-answer pairs for the competencies covered and for the objects referred.
-One question-answer pair might cover more than one competency or object, hence the total is higher than the size of the datasets.
-The Action and Scenery datasets have complementary strengths, with one focused more on action-justification competencies and one more on description and localisation.
-We created a collection of datasets for bringing language to autonomous driving.
-The total dataset size is 419.9k question-answer pairs, where a single data sample consists of a 4-second video clip at 1Hz.
-The total size of the dataset is about 10x larger than BDD-X [29], as shown in Table 3.
-Compared to prior datasets such as NuScenesQA [42], our dataset contains reasoning pairs in addition to object presence, description, and localisation.
-The answers are also more free-form and more complex, with an average answer length of 17.2 words versus 1.0 words in NuScenesQA.
-Examples of question answers pairs from LingoQA are shown in Appendix A.
-Our labeled autonomous driving training dataset consists of two complementary parts: the action dataset and the scenery dataset.
+Our labeled autonomous driving training dataset consists of two complementary parts: the action dataset and the scenery dataset. The breakdown of the datasets is shown in Table 3.
 Action dataset.
-This dataset was made from a recorded driving corpus of interesting events where the car’s behavior changes, such as decelerations, accelerations, lane changes, narrow gaps, and turns.
-Such events were succinctly labeled by driving operators with very short high-level descriptions of the situations and behavioral policies (e.g.
-“following lane, pedestrian on a zebra crossing, should stop”).
-Additionally, we added metadata for such events from various perception systems, such as traffic light presence, vehicles and pedestrian visual detectors, weather descriptors, as well as other metadata (speed, steering wheel position, and road type from the map data).
-Using this data, we developed prompt templates for (1) describing the current action and its justification and (2) a set of example questions and hints about what the answer should mention.
-Next, we used those prompts with GPT-3.5 to rephrase, answer, and extend the example questions using the provided action description and answer hints.
-We rebalanced events by bucketing by actions and behavioral policies and sampled up to 500 events from each bucket without replacement, leading to 24,577 video snippets with 167,774 question/answer pairs.
-Scenery dataset
-The scenery dataset was built to complement the action dataset by focusing on perception-related questions in addition to driving behaviours.
-The dataset was made by densely and thoroughly labelling three 30-minute driving sessions with the ELAN video annotation software[38].
-For the entire duration of the driving sessions, we provided short captions in about 15 different categories:
-  • Driver’s actions, and their justifications
-  • Driver’s attention
-  • Observations about relevant vehicles, pedestrians, and other road actors with their visual descriptions
-  • Observations about relevant road elements such as traffic lights, traffic islands, lane and intersection structures
-  • Miscellaneous observations about the environment, such as weather, tube stations, and buildings.
-Then, for every keyframe every second (1fps), we collect all annotations around this frame and build a textual description containing the driver’s actions and their justifications, the objects requiring the driver’s attention, and the observations.
-As opposed to the Action dataset, where recommended questions were provided to GPT-3.5 for rephasing, for the Scenery dataset, we asked GPT-4 to generate questions and answers using a set of generic prompts, but also using a prompt with the chain of thought specifically targeting perception questions.
-This forced GPT-4 to generate many diverse questions and answers.
-This led to a high-quality diverse dataset with about 43 QA-pairs per video.
+The action dataset was created from recorded driving sessions featuring notable events where the car’s behavior changes, such as decelerations, accelerations, lane changes, narrow gaps, and turns.
+These events were labeled by driving operators with concise, high-level descriptions of the situations and behaviors (e.g., "following lane, pedestrian on a zebra crossing, should stop").
+Additional metadata was included from various perception systems, covering traffic light presence, vehicle and pedestrian detection, weather conditions, and other data like speed, steering wheel position, and road type.
+Using this information, we developed prompt templates for (1) describing the current action and its justification and (2) generating example questions and hints for expected answers.
+We then used these prompts with GPT-3.5 to rephrase, answer, and extend the example questions based on the provided action descriptions and answer hints.
+Events were rebalanced by categorizing them based on actions and behavioral policies, and up to 500 events were sampled from each category, resulting in 24,577 video snippets with 167,774 question/answer pairs.
+Scenery dataset.
+The scenery dataset was designed to complement the action dataset by focusing on fine-grained perception-related questions.
+It was created by densely and thoroughly annotating three 30-minute driving sessions using the ELAN video annotation software.
+[40] During these sessions, brief captions were provided in approximately 15 different categories, including driver actions, justifications, attention, observations of vehicles, pedestrians, road elements (such as traffic lights, traffic islands and intersections), and environmental details (like weather and buildings).
+Annotations were collected every second (1 fps) for every frame to build textual descriptions that included driver actions, justifications, and observations.
+Unlike the Action dataset, where GPT-3.5 was used only for question rephrasing, GPT-4 was employed for the Scenery dataset to generate questions and answers.
+GPT-4 was prompted with chain of thought specifically targeting perception questions, resulting in a diverse and high-quality dataset with about 43 QA pairs per video.
+Dataset statistics.
 Our training dataset covers 9 different competencies: action (what the vehicle is doing), justification (why the action is taken), attention (what should be paid attention to in the current situation), identification (identifying an object given its description), localisation, description, counting, anticipation and reasoning given counterfactuals.
@@
 The complementary strengths of the datasets are apparent, with one focused on driving behaviours and one on perception tasks.
+Table 3: Dataset split.
+The Action dataset focuses on questions related to driving behaviours, the Scenery dataset focuses on perception capabilities, while the evaluation dataset is designed to probe a range of competencies.
+Scenarios QA pairs QA per scenario
+Action 24.5k 267.8k $\approx 10.9$
+Scenery 3.5k 152.5k $\approx 43.6$
+Eval. Dataset 100 1000 $10$
+Figure 2: Fine-tuning dataset statistics.
+The Action and Scenery datasets have complementary strengths, one focused on action-justification competencies and one on description and localisation.
+The number of question-answer (QA) pairs for the competencies covered and for the objects referred are shown.
+One QA pair might cover more than one competency or object, hence the total is higher than dataset size.
+Evaluation dataset.
+For evaluation we collected a small, low-density dataset from in-house human labelers, creating both the questions and the answers associated with the short videos.
+We labeled a small portion of held-out data on 500 human-generated questions using 20+ different evaluators to obtain our test set.
+Since answers are subjective and noisy, we labeled them twice, making sure the same evaluator does not receive the same question twice.
+After that, we manually reviewed the answers for semantic disagreements and mistakes.
+We relabeled such samples two more times and fixed the disagreements, preferring the semantics of the majority of responders but preserving maximal variety in the responses.
+Finally, we condensed this into 1k high-quality answers to 500 questions, with two correct but diverse answers per question, as shown in Table 3.
+The dataset evaluates a range of competencies, including action and justification, attention, description, localisation, identification, counting and anticipation, as shown in Appendix 0.A.
 4 Model Methodology
@@
 We train a model that consumes a short video segment and produces answers to autonomous driving-related questions.
-4.1 Architecture
-The LingoQA Baseline model architecture is based on recent VLMs [30, 35, 17] but enhances them by incorporating a video encoding strategy to process multiple frames from a video snippet, as shown in Figure 3.
 Figure 3: LingoQA Baseline model architecture.
@@
 The Q-Former outputs tokens and we feed the tokens from all frames along with chat history and questions into the LLM, which then predicts an answer.
+4.1 Architecture
+The LingoQA Baseline model architecture is based on recent VLMs [31, 35, 17] but enhances them by incorporating a video encoding strategy to process multiple frames from a video snippet, as shown in Figure 3.
 Vision encoder.
@@
 We use Vicuna v1.5 [13] with 7B parameters built on top of Llama-2 [47].
-The language model is auto-regressive and hence can be conditioned on textual inputs, as well as image tokens.
+The language model is autoregressive and hence can be conditioned on textual inputs, as well as image tokens.
 The training objective is to predict the next language token in a sequence.
@@
 In the second stage, we fine-tune the same parameters as in the previous stage, keeping the vision encoder frozen.
-Further details regarding training parameters are presented in Appendix F.
+Further details regarding training parameters are presented in Appendix 0.F.
+Stage 1: Pre-training for feature alignment.
+In the first stage, we pre-train the self-attention layers of the LLM and the vision encoder (QKV) on the GQA and SVIT datasets to align image features with the embedding space of the pretrained LLM.
+The GQA dataset [23] contains more than 22M questions over 113k images.
+The recently introduced SVIT dataset [61] contains 4.2M question-answer pairs over 108.1k images.
+We leverage initial weights from different models to accelerate the training process.
+We initialise the vision encoder using publicly available weights of OpenCLIP [24], the Q-Former from BLIP2 weights [32], and language model from Vicuna v1.5 [13].
+Stage 2: Fine-tuning for video QA.
+In the second stage, we fine-tune the model on our video question-answering Action and Scenery datasets described in Section 3.2.
+During the fine-tuning phase, each sample is composed of 5 frames taken from a 4-second span of video, accompanied by a QA-pair.
+The dataset used to fine-tune LingoQA Baseline is open-sourced.
+5 Empirical Evaluation on LingoQA
+Table 4: Empirical evaluation on LingoQA.
+Ablation study highlighting the impact of various modifications in training recipes, dataset composition, frame count, video processing techniques and language model.
 Ablation Lingo-Judge [%] $\uparrow$ BLEU $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$
@@
 Mistral-7B-Instruct [26] $58.00$ $13.80$ $18.33$ $64.21$
-Table 4: Empirical Evaluation on LingoQA.
-Ablation study highlighting the impact of various modifications in training recipes, dataset composition, frame count, video processing techniques and language model.
-Stage 1: Pre-training for feature alignment.
-In the first stage, we pre-train the model on the GQA and SVIT datasets to align image features with the embedding space of the pretrained LLM.
-The GQA dataset [23] contains more than 22M questions over 113k images.
-The recently introduced SVIT dataset [61] contains 4.2M question-answer pairs over 108.1k images.
-We leverage initial weights from different models to accelerate the training process.
-We initialise the vision encoder using publicly available weights of OpenCLIP [24], the Q-Former from BLIP2 weights [31], and language model from Vicuna v1.5 [13].
-Stage 2: Fine-tuning for video QA.
-In the second stage, we fine-tune the model on our video question-answering Action and Scenery datasets described in Section 3.3.
-During the fine-tuning phase, each sample is composed of 5 frames taken from a 4-second span of video, accompanied by a QA-pair.
-To facilitate further exploration of autonomous driving QA, we open-source the dataset used to fine-tune LingoQA Baseline.
-5 Empirical Evaluation on LingoQA
+5.1 Ablation Studies on LingoQA
 With the highly modular architecture of VLMs, the question remains what architectural components of the LingoQA Baseline model and dataset composition contribute the most to its performance?
 We conduct several ablation studies around the architecture and training paradigm described in Section 4.
-To this end, we investigate variations to the training strategy, training data composition, frame count, video fusion methods, and the use of different large language models.
+We investigate variations to the training strategy, training data composition, frame count, video fusion methods, and the use of different large language models, as shown in Table 5.
 The results are obtained by having each model generate one answer per question and then compare the predicted answer to the two ground truth answers.
-Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix G.
-5.1 Training Strategy
+Examples of comparisons between our baseline model’s answers and answers from other models from the ablations are presented in Appendix 0.G.
+Training Recipe and Dataset Mixture.
 The aim of the training strategy experiments is to understand how much the pre-training and the fine-tuning steps contribute to performance.
-It becomes apparent that fine-tuning leads to improved answers relevant to autonomous driving.
-The model fine-tuned on the LingoQA dataset has double the performance of the dataset that is pre-trained on generic VQA datasets.
-5.2 Training Datasets Mixture
-Table 4.2 shows the contributions of our two datasets, Action Dataset and the Scenery Dataset.
-Both datasets proved influential in improving model performance.
-We show that fine-tuning on the LingoQA dataset that we open source leads to a considerable improvement compared to general pre-training only.
-5.3 Impact of Frame Count
-We want to investigate the variation in VQA performance with decreasing and increasing the number of video frames fed into the model.
+Fine-tuning on the LingoQA dataset doubles the performance over generic VQA pre-training.
+The Action Dataset and the Scenery Dataset both prove influential in improving model performance.
+Impact of Frame Count.
+We investigated the variation in VQA performance with decreasing and increasing the number of video frames fed into the model.
 The base model contains 5 frames over a 4-second context.
-The performance declined when shifting from multi-frame video to a single image representation, which can be explained by the model not getting enough information to answer questions where temporal information is crucial.
-The performance when increasing the number of frames to 3, 5, and 7 remains relatively consistent.
-We hypothesize that this is due to the lack of effective video fusion, hence we highlight the opportunity for improved video encoding methods.
-5.4 Impact of Video Fusion Strategy
-Given how crucial the temporal context is for scenario understanding, this study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion.
+The performance declined when shifting from multi-frame video to a single image representation, while remaining close to the multi-frame baseline, showing that a certain proportion of autonomous driving scenarios can be solved from a single frame, as further discussed in Section 5.2.
+Nonetheless, to fully reach human-level multi-frame performance, video fusion is needed, as shown in Table 5.
+We conclude that both improved single-frame reasoning and video fusion are required.
+Impact of Video Fusion Strategy
+This study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion.
 The early-fusion method employs average pooling to condense features from the vision encoder prior to their incorporation into the Q-Former, producing a unified visual feature vector for language space projection.
-In contrast, the mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism.
-The late-fusion method, where individual frame embeddings from Q-Former output are fed into the LLM, allows it to resolve temporal relationships.
+The mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism.
+The late-fusion method feeds individual frame embeddings from the Q-Former output into the LLM.
 Our findings demonstrate that both mid-fusion and late-fusion are effective methods for incorporating video content into the model.
-Mid-fusion allows for a greater number of context tokens through the use of a predetermined number of video embeddings.
-Conversely, late-fusion shows a slightly enhanced performance by providing comprehensive frame information to the LLM.
-5.5 Impact of Large Language Model
+Impact of Large Language Model
 We investigate the impact that different Large Language Models have on the overall performance of our vision-language model.
-As shown in Table 4.2, the best score is achieved by Vicuna-1.5-7B [13] that our base model uses.
+The best score is achieved by Vicuna-1.5-7B [13].
 In the same family of models, Llama-2-7B [48] achieves comparable, but slightly lower performance.
-Mistral-7B [26], despite its promise of superior performance over Llama-2, proved less effective in our fine-tuning task.
-Notably, OPT-7B [60] demonstrates significantly lower performance, despite having a similar number of parameters.
-This discrepancy underscores the crucial role of the pretraining phase in the base language model’s effectiveness.
+Despite the promise of improved performance, Mistral-7B [26] is less effective in our fine-tuning task.
+OPT-7B [60] substantially underperforms the others, potentially due to the lower embedding size - 1048 compared to 4096 for all other models.
 5.2 Evaluation of SOTA Vision-Language Models
-To demonstrate the relevance of the newly proposed benchmark, we evaluate a series of SOTA vision-language models and compare them to human performance.
-The models shown in Table 5 are evaluated zero-shot.
-The best performing models achieve an accuracy as rated by humans of 56.67%, which is below human performance at 93.4%.
-This highlights the challenging nature of our proposed benchmark.
-Table 5: Evaluating vision-language models zero-shot. The performance of existing vision-language models is far from human capability, highlighting the challenging nature of the benchmark.
-Human Lingo-J GPT-4 BLEU METEOR CIDEr
-Human 93.4 96.6 4.68 81.04 52.92 361.77
-GPT-4V 56.67 59.6 3.19 6.30 12.35 42.82
-LLaVA 38.97 49.4 2.45 4.23 8.38 38.39
-FUYU 17.69 45.4 2.28 1.90 13.00 12.04
-6 Discussion
+To demonstrate the relevance of the newly proposed benchmark, we evaluate a series of SOTA vision-language models and compare them to human performance, as shown in Table 5.
+Human study.
+Human performance is evaluated on both video inputs and a single frame.
+We find a performance degradation from 96.6% to 81.8% without temporal context.
+The main failure modes include misclassifying parked cars as engaged in traffic, miscounting developing hazards such as motorcyclists and pedestrians over multiple frames, missing state transitions of traffic lights, and failing to predict the correct speed when behind a vehicle.
+These results indicate that while video understanding is required for reaching human multi-frame performance, improvements in single-frame reasoning are also crucial.
+Fine-tuned models.
+Our work identifies a notable 23% performance gap between single-frame LLaVA and single-frame human capability, underlining the benchmark’s significance for the autonomous driving and vision-language community.
+Our evaluation includes fine-tuning a single-frame LLaVA, a single-frame BLIP-2, and a text-only Vicuna-7B model on LingoQA, as shown in Table 5.
+Results reveal a performance lower bound of 38% for models lacking visual inputs.
+Notably, vision-language models, such as BLIP-2 and LLaVA, surpass this text-only baseline by 13% and 20.2%, respectively, with performance enhancements attributed to enhanced perceptual capabilities.
+Furthermore, LLaVA’s use of a larger CLIP crop size (336) compared to 224 improves performance.
+Zero-shot models.
+The best zero-shot model GPT-4V performance is still 37% below the multi-frame human performance.
+This highlights that further advancements are required for frontier models to fully solve the benchmark.
+The Lingo-Judge accuracy is impacted by the model response style, where long-form incorrect answers may receive high ratings, as it happens with FUYU.
+Further details are presented in Appendix 0.E.
+Table 5: Evaluating vision-language models on LingoQA. The performance of existing vision-language models is far from human capability.
+Category No. Frames Human Lingo-J BLEU METEOR CIDEr
+Human human study 5 93.3 96.6 81.04 52.92 361.77
+Human 1 - 81.8 10.64 15.01 64.45
+LingoQA fine-tuned models 5 57.1 60.8 15.00 18.56 65.62
+LingoQA 1 - 57.0 14.21 18.40 59.46
+LLaVA 1 - 59.0 12.5 18.5 57.0
+BLIP-2 1 - 52.2 13.0 17.4 60.1
+Vicuna-7B 0 - 38.8 10.1 15.2 51.0
+GPT-4V zero-shot models 5 56.61 59.6 6.30 12.35 42.82
+LingoQA 5 - 33.6 8.33 14.33 39.16
+LLaVA 1 38.97 49.4 4.23 8.38 38.39
+FUYU 1 17.69 45.4 1.90 13.00 12.04
+6 Discussion and Limitations
 Strengths of Lingo-Judge.
 The strength of our contribution comprises proposing a classifier that is highly correlated with human inputs and efficient to run.
-In conjunction with the evaluation dataset that we propose, it becomes a useful tool for benchmarking vision-language models for autonomous driving on the video question answering task, which has been historically challenging to evaluate in a consistent fashion.
+In conjunction with the evaluation dataset that we propose, it becomes a useful tool for benchmarking vision-language models for autonomous driving for visual question answering task, which has been historically challenging to evaluate in a consistent fashion.
 With this contribution, autonomous driving research can be accelerated by providing a reliable, efficient, and easy-to-interpret benchmark.
 Limitations of Lingo-Judge.
-This work still has a few limitations that we discuss and provide guidelines as to how their effects may be mitigated.
-The classifier has been trained for autonomous driving evaluation and has been shown effective for this purpose.
-Hence, the classifier can be used for evaluation when paired with the benchmark dataset that we release with the paper.
+The Lingo-Judge is specifically tailored for open-vocabulary evaluation on the LingoQA benchmark.
+While it demonstrates high speed and accuracy compared to larger models like GPT-4, it is designed akin to the model described in the TruthfulQA [34] paper, which suggests that such specialized models are not expected to generalize well to new questions.
 Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset.
 The same response style is adopted in the LingoQA training sets and the models.
-Further details regarding generalisation to response styles is studied in Appendix E.
+Further details regarding generalisation to response styles is studied in Appendix 0.E.
 Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer.
@@
 7 Conclusion
-In this paper, we introduced a novel benchmark for Video Question Answering for autonomous driving.
-The benchmark consists of a evaluation dataset, learned classifier-based metric Lingo-Judge that is highly correlated with human evaluation, a comprehensive high-quality training dataset for autonomous driving.
-The fast feedback from employing Lingo-Judge facilitates effective exploration in the video QA field.
+In this paper, we introduced a novel benchmark for Visual Question Answering for autonomous driving.
+The benchmark consists of an evaluation dataset, learned classifier-based metric Lingo-Judge that is highly correlated with human evaluation, a comprehensive high-quality training dataset for autonomous driving.
+The fast feedback from employing Lingo-Judge facilitates effective exploration in the visual QA field.
 Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of end-to-end autonomous driving systems.
@@
 This work was possible through the help of many colleagues across Wayve.
-We would like to thank Elahe Arani for the valuable discussions and feedback on the report.
 In particular we would like to acknowledge the support from: Anthony Hu, Miguel Sanchez Lohff, Lorenzo Bertoni, Charlie Lyons-Rothbart, Emma Wang, Harriett-Rose Follas, Kyle Esecson, Ben Foxall, Naama Zahavi, Ruben Diaz, Rudi Rankin, Tilly Pielichaty, Sarah Belghiti, Giulio D’Ippolito, Dan Reisman, Alex Persin, Fergal Cotter, Przemyslaw Mazur, Thomas Sajot, Giacomo Gallino, Alex Garcia Mayans, Tim Geypens, Robin Tweedie, Rebecca Hills.
-Appendix A LingoQA Dataset Examples
+Appendix 0.A LingoQA Dataset Examples
 Further examples on the capabilties existent in the training and the evaluation datasets are shown in Figure 4.
@@
 Further examples from the evaluation benchmark are also included in the overview video accompanying the submission.
-Notably, any personal identifiable information, such as faces and plate ID’s, has been annonymised in the dataset.
-Figure 4: LingoQA dataset examples. From left to right: scenery dataset, action dataset, and evaluation dataset.
-Appendix B Lingo-Judge Examples
+The dataset statistics for the evaluation dataset are shown in Figure 5.
+Notably, any personal identifiable information, such as faces and plate ID’s, has been anonnymised in the dataset.
+Figure 4: LingoQA dataset examples.
+From left to right: scenery dataset, action dataset, and evaluation dataset.
+Further video examples are provided in the supplementary material accompanying the submission.
+Figure 5: Evaluation Dataset Statistics. The evaluation dataset assesses a range of competencies and includes a wide range of objects relevant for autonomous driving.
+Appendix 0.B Lingo-Judge Examples
 We present additional qualitative examples from our evaluation dataset in Table 6, alongside predictions from our base model and corresponding metrics for each individual sample.
@@
 Lingo-Judge correctly identifies that the statements described by the model are false.
-Question
-Label
-Prediction
-CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
-How many pedestrians are crossing the road?
-Zero pedestrians
-There are no pedestrians crossing the road.
-23.18 5 5 0.96 True
-What is the road speed limit?
-20 mph - it is written on the road
-The road speed limit is 20 mph.
-97.91 4 3 0.95 True
-How many cars are driving in your direction?
-None
-There are no cars driving in my direction.
-0.11 4 3 0.96 True
-Which vehicle should you follow if any?
-The motorcyclist.
-If any, I should follow the motorcycle ahead.
-0.42 4 5 0.95 True
-What is the current action and its justification? Answer in the form “action, justification”
-Slow down, there is a stationary van infront of us
-I am decelerating because of the stationary truck ahead.
-34.79 5 5 0.96 True
-What is the current action and its justification? Answer in the form “action, justification”
-Stop, Red light
-I am stopping because the traffic lights to go straight are red.
-18.62 5 5 0.95 True
-How many cyclists can you see?
-I can see 3 cyclists
-I can see two cyclists.
-150.29 1 2 0.05 False
-What color are the traffic lights showing?
-The traffic lights are showing green
-The traffic lights are showing red.
-329.36 0 1 0.05 False
-What action are you taking with respect to the cyclist?
-Overtaking them on the right and keeping the speed
-I am overtaking the cyclist on the left.
-349.52 2 2 0.10 False
-In which direction is the bus driving?
-The bus is driving in the opposite direction
-The bus is driving in the oncoming direction.
-404.65 4 5 0.31 False
-Are there any parked car on the side of the road?
-Yes, there are two cars parked on the right of the road
-No, there are no parked cars on either side of the road.
-142.40 0 0 0.05 False
-Is acceleration necessary in this situation? If so, provide the reason.
-No. We should decelerate in this situation because there is a vehicle stopping ahead of us.
-No, acceleration is not necessary in this situation as I am already driving at the speed limit.
-177.96 3 3 0.31 False
 Table 6: Qualitative comparison of metrics.
@@
 Green color indicates agreement with human judgement, while orange and red show disagreement.
-Question
-Label
-Prediction
-CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
-Is any vehicle stopped at the zebra crossing?
-There is no zebra crossing.
-No, there are no vehicles stopped at the zebra crossing.
-106.19 4 2 0.93 True
-What elements are capturing your attention and what is the reason behind it?
-I am making sure there is no vehicle on the left lane I am making my lane change to. I am also looking at the traffic light to make sure it stays green while I am crossing the intersection.
-I am paying attention to the bus lane on the left because I need to maintain a safe distance from it.
-0.91 3 1 0.57 True
-What is the current action and its justification? Answer in the form ”action, justification”
-We are taking a left turn, because we are at a t-junction. Then we keep a safe distance from the cyclists in front.
-I am decelerating to keep a safe distance from the cyclist ahead of me.
-46.45 2 2 0.32 False
+Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
+How many pedestrians are crossing the road? Zero pedestrians There are no pedestrians crossing the road. 23.18 5 5 0.96 True
+What is the road speed limit? 20 mph - it is written on the road The road speed limit is 20 mph. 97.91 4 3 0.95 True
+How many cars are driving in your direction? None There are no cars driving in my direction. 0.11 4 3 0.96 True
+Which vehicle should you follow if any? The motorcyclist. If any, I should follow the motorcycle ahead. 0.42 4 5 0.95 True
+What is the current action and its justification?
+Answer in the form “action, justification” Slow down, there is a stationary van infront of us I am decelerating because of the stationary truck ahead.
+34.79 5 5 0.96 True
+What is the current action and its justification? Answer in the form “action, justification” Stop, Red light I am stopping because the traffic lights to go straight are red. 18.62 5 5 0.95 True
+How many cyclists can you see? I can see 3 cyclists I can see two cyclists. 150.29 1 2 0.05 False
+What color are the traffic lights showing? The traffic lights are showing green The traffic lights are showing red. 329.36 0 1 0.05 False
+What action are you taking with respect to the cyclist? Overtaking them on the right and keeping the speed I am overtaking the cyclist on the left. 349.52 2 2 0.10 False
+In which direction is the bus driving? The bus is driving in the opposite direction The bus is driving in the oncoming direction. 404.65 4 5 0.31 False
+Are there any parked car on the side of the road? Yes, there are two cars parked on the right of the road No, there are no parked cars on either side of the road. 142.40 0 0 0.05 False
+Is acceleration necessary in this situation?
+If so, provide the reason.
+No.
+We should decelerate in this situation because there is a vehicle stopping ahead of us.
+No, acceleration is not necessary in this situation as I am already driving at the speed limit.
+177.96 3 3 0.31 False
 Table 7: Failure Cases of Lingo-Judge.
@@
 Green color indicates agreement with human judgement, while orange and red show disagreement.
+Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
+Is any vehicle stopped at the zebra crossing? There is no zebra crossing. No, there are no vehicles stopped at the zebra crossing. 106.19 4 2 0.93 True
+What elements are capturing your attention and what is the reason behind it?
+I am making sure there is no vehicle on the left lane I am making my lane change to.
+I am also looking at the traffic light to make sure it stays green while I am crossing the intersection.
+I am paying attention to the bus lane on the left because I need to maintain a safe distance from it.
+0.91 3 1 0.57 True
+What is the current action and its justification?
+Answer in the form "action, justification" We are taking a left turn, because we are at a t-junction.
+Then we keep a safe distance from the cyclists in front.
+I am decelerating to keep a safe distance from the cyclist ahead of me.
+46.45 2 2 0.32 False
 Figure 6: Classifier examples. Examples of Lingo-Judge outputs compared to GPT-4.
-Appendix C GPT-4 Grading
+Appendix 0.C GPT-4 Grading
 In this section we provide an overview of the implementation details for the evaluation method using GPT-4 with and without Chain-of-Thought (CoT) [52] prompting.
@@
 GPT-4 is provided with a prompt that contains the question, the ground truth answers, and the model response, and is requested to directly provide a grade for the student.
-Appendix D Lingo-Judge Correlation Study
+Appendix 0.D Lingo-Judge Correlation Study
 We show that Lingo-Judge exhibits a higher correlation to human judgment than commonly-used language-based metrics, and than GPT-4.
@@
 We note that not only does Lingo-Judge provide higher correlation, it also provides tighter confidence intervals than the other metrics.
-Appendix E Lingo-Judge Generalisation
+Appendix 0.E Lingo-Judge Generalisation
 To investigate the generalisation abilities of the Lingo-Judge, we examine the performance of the model on a range of answer styles.
 In particular, we evaluate vision-language models with varying architectures namely GPT4-V, LLaVA and FUYU.
-Table 8 shows the performance of the Lingo-Judge as measured by the validation accuracy.
+We also employ human labelling to obtain ground truth performance of these models.
+Table 8 shows the performance of the Lingo-Judge as measured by the validation accuracy, as well as human, Lingo-Judge and GPT-4 ratingd for comparison.
+The Lingo-Judge rates FUYU at 45.4%, compared to 17.69% as rated by humans.
+However, the Lingo-Judge rates GPT-4V concise at 56.6% compared to 56.67% by humans.
 This highlights that the model performs the best on short, concise answers, akin to those of humans.
-Table 8: Robustness to response styles. Investigation into the impact of the response style on validation accuracy.
-Lingo-J Val. Acc.
-LingoQA 89.50
+Table 8: Robustness to response styles.
+Investigation into the impact of the response style on validation accuracy.
+The Lingo-Judge accuracy is limited on mostly incorrect but long-form answers, with the main failure mode being that these models are rated higher than they should be, compared to the human evaluation.
+Human Lingo-Judge GPT-4 Lingo-J Val. Acc.
+LingoQA 57.10 60.8 3.30 89.50
 GPT-4V
-few-shot (FS) 83.27
-concise (C) 81.63
-unprompted (U) 83.06
-incorrect (I) 89.59
+few-shot (FS) 57.69 64.6 3.39 83.27
+concise (C) 56.67 59.6 3.19 81.63
+unprompted (U) 54.61 59.6 3.24 83.06
+incorrect (I) 17.69 26.2 1.55 89.59
 LLaVA
-concise (C) 81.43
-unprompted (U) 78.12
-FUYU 64.89
+concise (C) 38.97 49.4 2.45 81.43
+unprompted (U) 28.28 41.8 2.69 78.12
+FUYU 17.69 45.4 2.28 64.89
+Table 9: Correlation study metrics.
+Metrics from different models on our evaluation dataset used in the correlation study in Table 1.
+For reference, we also present metrics for answers provided by human labellers.
+“Human” is the average of inference output scores in range $\left[0,1\right]$ where 0 is worst and 1 is best, as described in section 3.1.
 Lingo-Judge [%] $\uparrow$ BLEU $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$ GPT-4 $\uparrow$ Human $\uparrow$
@@
 Human labellers group B $91.2$ $61.72$ $42.57$ $267.87$ $4.3$ $0.894$
-Table 9: Correlation study metrics.
-Metrics from different models on our evaluation dataset used in the correlation study in Table 1.
-For reference, we also present metrics for answers provided by human labellers.
-“Human” is the average of inference output scores in range $\left[0,1\right]$ where 0 is worst and 1 is best, as described in section 3.2.
 Figure 9: Correlation trends. Correlation trends of the average grade of models compared to the average human-grades, for different metrics.
 Figure 10: Correlation coefficients. Correlation coefficients of the average grade of different models vs. the average human-grades, for different metrics.
-Appendix F Training Parameters
+Appendix 0.F Training Parameters
 In this sections we present further details on the training parameters used for the LingoQA Baseline.
@@
 The overall training time was  20h for pre-training and  5h for fine-tuning on an NVIDIA A100 8GPU 80GB machine.
+Table 10: Training parameters. This table shows the training parameters utilised for the pre-training and for the fine-tuning stages respectively.
 Parameter Pre-training Fine-tuning
@@
 Weight decay 0.1 0.1
-Table 10: Training parameters. This table shows the training parameters utilised for the pre-training and for the fine-tuning stages respectively.
-Appendix G LingoQA Baseline Examples
+Appendix 0.G LingoQA Baseline Examples
 Figure 11: Examples of model outputs on the LingoQA benchmark.